A Bibliography of Publications of Stanisław M. Ulam

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

05 June 2025 Version 2.66

Abstract This bibliography records publications of Stanisław Ulam (1909–1984).

Title word cross-reference

$17.50 [Bir77]. $49.95 [Bár04]. $5.00 [GM61]. α [OVPL15].

-Fermi [OVPL15].

150 [GM61]. 1949 [Ano51]. 1961 [Ano62]. 1963 [UdvB^{+}64]. 1970 [CFK71]. 1971 [Ula71b]. 1974 [Hua76]. 1977 [Kar77]. 1979 [Bud79]. 1984 [DKU85].

2000 [Gle02]. 20th [Cip00]. 25th [Ano05].

49.95 [Bár04].

60-year-old [Ano15].

7342-438 [Ula71b]. ’79 [Bud79].

Abbildungen [MU31, SU34b, SU34a, SU35a, Ula34, Ula33a]. Abelian [MU30]. Abelsche [MU30]. Above [Mar87]. abstract [Ula30c, Ula39b, Ula78d]. abstrakten [Ula30c]. accelerated [UW64]. Accelerates [Gon96b]. Adam [Gle02]. Adaptive [Hol62]. additive [BU42, Ula30c]. Advances [Ano62, Ula78d]. Adventures [Met76, Ula76a, Ula87d, Ula91a, Bir77, Wil76]. Aerospace [UdvB^{+}64]. ago [PZHC09]. Alamos [DKU85, MOR76a, HPR14, BU90, Ula91b]. Albert [RR82]. algebra [BU78, CU44, EU45b, EU46, Gar01a, TWHM81]. Algebraic [Bud79, SU73]. Algebras [BU75, BU76, EU50d]. Algorithm [JML13]. Algorithms [Cip00]. allgemeinen [Ula30d]. America [FB69]. American [UdvB^{+}64, Gon96a, TWHM81]. Analogies [BU90]. analogy [Ula81c, Ula86]. Analysis [Goa87, JML13, RB12, Ula49, Ula64a, Jun01]. Andrzej [Ula47b]. anecdotal [Ula82c]. Annual [Ano05]. any [OU38a, Ula38b]. applicability [Ula69b]. application [EU50d, LU34]. Applications [Ula56b, Ula56a, Ula81a]. Applied [GM61, MOR76b, TWHM81, Ula67b]. appreciation [Gol99]. Approach [BPP18, Ula42]. Approximate [HU42, HU47, HU41, HU45]. Approximately [HU51, HU52]. April [Fla84, Gle02]. area [Ula43c]. Arithmetic [Bud79]. arithmetical [MU53, MUxx]. Arlington [TWHM81]. arms [McM05]. Arthur [Ula75]. Astronautics [Ula58f]. astronomy [FF91]. astrophysics [RR82]. Atomic [Ula60b, Dys02]. Attempts [Ula87c, US71, US86]. Attractors [Ste87, SZ10]. Autobiography [Gam70b, Gam70a]. Automata [Bur70, vNB66]. automorphism [SU33b, SU37]. Automorphismen [SU33b, SU37]. aux [UK33]. Away [HM90].

Ban [Ula63a]. Banach [Ula46, Ula78a]. Barrier [Rot86a, Rot85, Rot89]. basic [LU33b]. before [Ula58f, Ula60b, Ula63a, Ula71b]. Beginning [Met87]. behavior [Ula56c]. Bemerkung [SU35a]. Bemerkungen [SU34b]. beneath [Wil76]. Berlin [Bud79]. Berlin/Wendisch [Bud79]. Berlin/Wendisch-Rietz [Bud79]. Bernstein [Ula29b]. Best [Cip00]. bete [CU^{+}66]. Betrag [Ula30a]. Bets [HPR14]. between [EU50d, BU90]. Beyond [Ano05]. bibliography [KW64]. binary [SU59, Ula68a]. Biographies [Ula82a]. Biology [Ano87d, BSSU72, Ula78d, Wat06]. Biomathematics [Myc87, Ula72b]. Birth [PZHC09, McM05]. body [Ula62a]. Bomb [HM90, Sul84, Zor24, BHM91, For15, HM97, Kno90, WG17]. Book [Asp88, Bár04, Bir77, GM61, May69, Met76, Moo82, Str98, Gar01a, Mau81, Ula57d, Mau15, Ula77, Ula82c]. Books [Wil76, Ula58a]. Borsuk [Bár04]. bounded [Ula38a]. Brain [URR86c, Ula87c, Ula58d]. branching [Ula68a]. Brilliance [FE85]. Brown [Ula38f]. Building [For15]. Burning [EU50c, EU50b].

C [Asp88]. Cafe [Rot87, Mau81, Mys90, Rot90, Mau15, Ula69a, Ula69c]. calcul [LU34]. calculate [Fro99b]. Calculating [Ano51]. calculation [LU33b]. Calculations [Ula61b, Ula62d, Ula13]. Calculus [BU75, BU78]. caractéristique [AMU35]. Cardinals [Arr89, Myc87, Rid89, CES89]. Carlo [Asp88, Ben16, BHMU85, Dys12a, Eck87, HPR14, Hur85, JML13, KW64, MU49b, Ula56b, Met87, Ula48, Ula51, Ula56a, Ula61b, Ula62d, Ula13]. case [Ula71b]. Categorial [Bud79]. category [OU38a, Ula31]. cathedral [Dys12b]. Cellular [Bur70]. Center [Ano05]. centre [UW64]. century [Har06, Kur80, Moo82, Cip00, Har10, MHR80]. certain [GLMU56, SU55, SU59, Ula33b, Ula33a, Ula63c, Ula68a, Ula85]. Chain [JML13]. challenge [Ula75]. chance [Ula75]. changed [Har06]. changing [Ula65b]. Chaos [Arr89, Rid89, CES89]. characteristic [AMU35]. charakteristische [AMU31]. Chess [KSU^{+}57, SU57a, SU57b]. Chirikov [FS10]. circle [SU35a]. City [UdvB^{+}64, CFK71]. classes [BU42]. classic [Gar01a]. closer [Har10]. cluster [UW64]. coefficient [UK33]. Collaboration [Ber10]. collaborators [BU90]. Collapse [Ula66a, UW64]. Collection [GM61, MHR80, Ula60a]. colossal [Gar01a]. combinations [UvN47]. combinatoire [UK32]. combinatoires [LU34]. Combinatorial [Ula49, Ula64a, EU44a, HU60, LU34, SU55, UK32, Ula72a]. Combinatorics [Bár04]. Comments [BHMU85]. Committee [Ula60b, Ula63a]. Communication [Ula63a]. compact [SU35b]. Complete [Gol99, Ula30c, Ula38b]. completely [BU42]. complexity [BSU71, Ula81c, Ula86]. Complicated [Pat87]. Composition [BU77b]. Computation [Ano51, BU77a, Bud79, Kar77, LU34]. Computational [BU77b, Wat06]. Computations [BHMU85, Hur85, Ula68a, Ula56d, Ula62a]. Computer [BSU69, Mac92, Ben16]. Computers [Asp88, Ula64b, URR86c, SU63, SU64, Ula58c, Ula58d, Ula62c, URR86d]. Computing [MHR80, Ula80d, URR86e, SU55, SU57a, SU57b, UP53, Ula57c, Ula72a, Ula80c]. concept [Ula32]. conceptual [Ula78d]. Concerning [Ula29a, US71, US86]. Conditions [BLP09]. Conference [Ano62, Ano05, Bud79, DKU85, Kar77, CFK71, TWHM81]. conjecture [Myc77]. Connected [Ula62b, SU35b]. connexe [SU35b]. Considerations [EU50a]. constant [SU34b]. construction [Ula67c]. constructive [Ula81c, Ula86]. contents [Ano87c]. continues [PZHC09, SU33a, UK33]. continuous [HU42, HU47, SU33a, UK33, Ula34]. contribution [RW12, Ula30a]. Contributions [BSW85, Goa87]. controversy [BHM91]. Convergence [JML13, Bar86]. Conversations [Ula87a]. convex [HU51, HU52]. Copernicus [Ney74]. Corporation [Ula71b]. correspondence [EU50d]. Cosmology [Rei72]. Court [Ula71b]. Cover [Ano87c]. covering [KW64]. Creating [Zor24]. creatrice [Ula66c]. cube [Myc77]. curves [Ula43c].

D [EU50b]. dans [LU34]. Debray [Ula75]. Debray-Ritzen [Ula75]. decompositions [CEG^{+}79, Ula33b]. dedicated [MOR76a, MOR76b, RR82, Rot86b]. défi [Ula75]. defined [GLMU56, SU67a]. definitions [Ula65b]. Delivered [Ano62]. Demons [Dys12c]. dense [UW64]. d’ensembles [UK33]. densities [Bar86]. dependent [BSU69]. Designer [Zor24]. d’espaces [MU32]. determination [EU48b]. deterministic [UvN47]. Deterrence [Mac92]. Detonations [TU51]. Deuterium [EU50c, EU50b]. Development [Ula60b]. devices [Ula66d]. didn’t [MHM90]. difference [Ras10]. differential [UH54]. diffusion [RUvN47, Ula48, Ula61c]. Digital [Ano51, Dys12b]. dilettante [Ula63b]. Dimensions [SU60]. discoveries [Ula68c]. discovery [PZHC09]. Discrete [Has87]. discrimination [Ula82e]. discussion [UdvB^{+}64]. Display [SUW64]. dissertation [Ula97]. distance [MU69]. distributed [MU49a]. Distribution [SUW64, SU67b, Ula38b]. District [Ula71b]. Division [UdvB^{+}64]. doctoral [Ula97]. Dr. [Goo04]. d’un [AU35, SU35b]. during [Ula38f]. dynamical [Fro99b, RW12, SZ10, Ula56c, Ula61a, BC82]. Dynamics [BCP13, BPP18, Wei97, Str98].

e-Abbildungen [Ula33a]. e-isomorphic [Ula39a]. e-transformations [Ula33a]. Early [BHMU85, Har87, Hur85]. Eclectic [FE85]. Editors [Cip00]. education [Ula67b]. Edward [Goo04, Har10]. Effects [AMS19, BP11]. Effort [Gon96a]. Eigenschaft [AMU31]. eigenvalue [AMU31]. Eindeutigkeit [Ula30b]. Einige [Ula31]. Einstein [RR82]. Electronic [Ula62c, Ben16, SU57a, SU57b, SU63, SU64, Ula56d]. elementary [US71, US86]. ellipsoid [AMU35]. ellipsoïde [AMU35]. Ellipsoides [AMU31]. ellipsoids [AMU31]. endomorphisms [EU48b]. Energy [Ula60b, Ula58e, UdvB^{+}64, Ula60b]. Engineer [Bec62, Bec13]. ENIAC [HPR14]. Enrico [Gol99, PZHC09]. Entretien [Ula75]. entropy [EU69, Fro99b]. equation [EU50e]. Equations [RB12, EU68, Jun01, Ras10, Ula48]. Equipartition [BP11]. equivalence [BU42, OU38a, Ula43b]. Ergodic [Pat87, vNU45, Ula38e, Ula56c]. Ergodicity [Myc87]. ergodiques [Ula38e]. espaces [LU34]. Esquisse [Ula87b]. Essays [Bur70, MHR80, MOR76b, MOR76a]. Euclidean [SU33a, SU34a]. euclidiennes [SU33a]. euklidischen [SU34a]. Europe [FB69]. evolution [Ula66b, Ula67a]. Evolutionary [BSSU74, US71, US86]. Excerpts [Ula87d]. Existence [OU38b, OU39]. experiment [DPR05, PZHC09]. Experimental [PZHC09]. experimentally [Ula72a]. Experiments [KSU^{+}57, SU57a, SU57b, SU55]. Exploration [Ula58f]. explosions [EU55]. Exponents [BPP18]. expressions [UH54]. external [EU55]. extracting [Ula58e].

F [Hua76]. F.U.N. [Ras10]. fallout [BHM91, Han97]. FCT [Bud79, Kar77]. FCT-Conference [Kar77]. Fe [Ano05]. Feigenbaum [Ano93]. Fermat [EU71]. Fermi [Ano05, OVPL15, Str98, AMS19, Ano15, BLP09, BP11, BCP13, BPP18, BI04, BI05, DPR05, Dau08, Gol99, PZHC09, Ula55, Wei97]. Fifty [BI05]. Figures [SU60, Ula62b]. Finite [BP11]. Finite-Size [BP11].

First [Zor24, Ben16, CFK71, MOR76a, OU38a]. five [Har06]. fixed [Ula30e, Ula47a]. Fixpunktes [Ula30e]. Florida [BC82]. flows [Ula82f]. Fluids [Has87]. Foreign [Ula63a]. Foreword [Ula70, Ula80b]. Formalized [Ula47b]. formulate [Ula66b, Ula67a]. formulations [Ula67c]. foundations [Ula85]. fourth [TWHM81]. frameworks [Ula85]. Fray [Mar87]. French [AU35, AMU35, LU34, SU33a, SU35b, US31, UK32, UK33, US33, Ula38e, Ula75]. Frontiers [Ula60b]. Fuchs [Ber10, MHM90]. function [MU53, MUxx, Ula47a]. Functional [Ras10, RB12, Jun01]. functions [HU42, HU47, HU51, HU52, Ula29a, Ula30c, Ula38f, Ula43b, Ula59a, Ula59b]. Fundamentals [Bud79, Kar77]. Funkcje [Ula59a, Ula59b]. fusion [Rei72]. Future [Ula58c, Ula81a, UdvB^{+}64].

galaxy [UW64]. Game [Ano62, Mac92, Gar01a]. games [MU87b, Ula56b, Ula56a, Ula74b]. Gamow [Rei72, Ula72c, URR86a]. Garwin [Zor24]. Gave [HM90]. GDR [Bud79]. genealogical [MU69]. general [BU42, Ula30d, Ula38b, Ula43c, Ula67c, Ula97]. Generalizations [Ula71a]. generalized [Ula29b, Ula80a]. generalizing [EU66]. générateurs [AU35, SU35b]. Generators [BU76, AU35, SU35b]. Genesis [Wei97, Str98]. Genetics [Goa87]. Geneva [Yor87]. Genève [CU^{+}66]. Genius [Mac92]. Geometrical [Ula42, SU67a]. Geometry [Bár04, Gar01a]. George [Rei72]. Gerard [Ula30b]. Gerardenmengen [Ula30b]. German [AMU31, LU33a, LU33b, MU30, MU31, SU33b, SU34b, SU34a, SU35a, SU37, Ula30a, Ula30b, Ula30d, Ula30e, Ula30c, Ula31, UB31, Ula32, Ula33b, Ula33a, Ula34]. Gesetze [LU33a]. gewisse [Ula33b, Ula33a]. Gian [Asp88]. Gian-Carlo [Asp88]. Google [SZ10]. graph [CFK71]. Graphs [Gra87, CEG^{+}79]. gravitational [Ula58e]. grossen [LU33a]. Group [vNU47, AU35, EU66, SU35a, SU35b, US33]. groupe [AU35, SU35b, US33]. groups [EU44b, EU45a, MU30, SU33b, SU37, Ula42]. Growth [SU60, Ula62b, HU60, SU67a]. Grundlagen [LU33b]. Gruppe [SU35a]. Gruppen [MU30].

H [BHM91, For15, HM90, HM97, Kno90]. H-Bomb [HM90, BHM91, HM97, Kno90]. half [Kur80, Moo82]. Hamilton [EU50e]. Harvard [Ano51]. hazard [Ula75]. Hearings [Ula63a, Ula58f, Ula60b]. held [Bud79, CFK71, DKU85, TWHM81]. heritage [Ney74, TWHM81]. Heterocatalytic [TU51]. Heuristic [PU59, UP53]. high [UP53]. Hilbert [Myc77]. Hiroshima [Yor87]. History [MHR80, BSU69, DPR05, For15, RR82, RW12, Ula82c]. history-dependent [BSU69]. Homage [Ula55]. Homeomorphisms [vNU47, OU41]. homme [CU^{+}66]. Honeywell [Ula71b]. honor [Hua76]. Hotel [Ano05]. House [Ula58f]. hull [BU68]. hulls [MU67]. Hydrodynamic [TU51]. Hydrodynamics [Har87, PU59]. Hydrogen [Sul84, Zor24, WG17]. Hyers [Jun01]. Hypothesis [Pat87].

Ideas [Ula73, Ula72b, Ula78d]. Ignition [EU50c, EU50b]. II [BC82, Ula31, EU49b, EU50b, Ula43d, UE48b, Ula59b]. II-er [Ula31]. III [EU49c, Ula44, UE48c]. Impact [Ano05]. implications [Ula68c]. Incorporated [Ula71b]. independantes [LU34]. inequalities [Ras10]. Infamous [Wel12]. Infinite [Ula57a, Ula64a, MU30, MU49a]. Infinities [Ula74a]. infinity [Gar01a]. influential [Har10]. Informal [Gam70b, Gam70a]. Initial [BLP09]. inni [Ula78a]. inspire [PZHC09]. Instabilities [BHJ^{+}87]. Institute [Hua76]. Integer [BU79]. Integer-Valued [BU79]. integers [GLMU56]. Integrable [BCP13, BPP18]. intellectual [FB69]. interacting [EU69]. Interaction [Ula80d, URR86e]. International [Ano05, Kar77, BC82]. Internationales [CU^{+}66]. Interscience [GM61]. interval [Fro99a]. Interview [Bar81, Ula75, Ano82, Ano93, Fei85]. Interviews [AA85, AA08]. Introduction [Ula82d, BSSU72]. invariance [Myc77]. invariant [OU39]. Invarianten [Ula33a]. invariants [Fro99b, Ula33a, Ula38d]. investigated [Ula56d]. involving [Ula48]. isometric [MU31]. isometries [HU41, HU42, HU45, HU47]. isometriques [MU32]. isometrische [MU31]. isomorphic [CEG^{+}79, Ula39a]. isomorphism [RW12]. isomorphisms [Ula71a]. isoperimetric [Ula43c]. Iteration [Ste87, Bar86]. iterations [Ula82f]. Itself [Ula87c, SU35a]. ivory [Wil76].

J [Bár04]. J. [McM05, PC06]. Jacobi [EU50e]. John [Ber10, Eck87, Mac92, PZHC09, Réd05, Ula58b, Ula58d, UKTS68, UKTS69, Ula82a, Ula82b, URR86b, URR86c]. Joint [Ula60b, Ula60b]. Judging [Har10]. June [CFK71].

Kac [Kem70, May69, Ano82, Ano93, Fei85, Rot86b]. Kategorie [Ula31]. Kawiarni [Ula69c]. Kazimierz [Moo82, Ula81b, Ula83]. kinematics [CU44]. Klaus [Ber10]. know [MHM90]. Koestler [Ula75]. Kornik [Kar77]. Kreislinie [SU35a]. Ksiegi [Ula78b]. Kuratowski [Moo82, Ula81b, Ula83].

Laboratory [Ano51, DKU85]. lady [Dau08]. Large [Ano51, EU50c, EU50b, LU33a]. Large-Scale [Ano51]. last [EU71]. late [Ula58d]. Laws [BLP09, LU33a]. Learning [Myc87]. Lebesgue [US31]. Lectures [Bár04, SU73, Ula38f]. Legacy [Ano87d, Ano87e, Ano87f, Ano05, Arr89, Rid89, CES89]. length [Ula43c]. Lenses [TU51]. Letter [Han92, Han97, HM97, Kno90, Zie77]. Letters [BHM91, MHM90, Réd05]. lié [UK33]. Life [Arr89, Rid89, CES89, PC06]. Limit [BP11]. Line [Gam70b, Gam70a, MU49a, SU35a]. Linear [JML13, SU63, SU64, Ula63c]. Logic [Kem70, May69, KU68, KU79, KU92]. Logical [Hol62]. logika [KU71]. logique [KU73]. look [Har10]. looks [Ula80c]. Lorentz [EU66]. Lost [Mys90, Rot87, Rot90, Mys90, Rot90]. Lyapunov [BPP18].

M [Ano87a, Bar85, Bar08, BSW85, Bir77, GM61, Kem70, Met76, MOR76a, MOR76b, Rot84, BU90, Ula94, Wil76]. M. [US31]. machine [Ula66c]. Machinery [Ano51]. machines [SU55, SU57a, SU57b, UP53, Ula56d, Ula57c, Ula72a]. Making [Yor87]. Manhattan [Ben16]. manifolds [Ula34]. Mannigfaltigkeiten [Ula34]. many [Ula62a]. many-body [Ula62a]. map [Bar86, FS10]. Maps [Ste87, Fro99a, Ula34]. March [Gle02]. Marcin [Fla84, Myc90]. Marian [Ula57b]. Mark [Asp88, May69, Ano82, Ano93, Fei85, Kem70, Rot86b]. Markov [JML13]. Martians [Har06]. Mass [EU50c, EU50b, MU49a]. Massachusetts [Hua76]. Massbegriff [Ula32]. Masses [Ula30b]. Massfunktionen [Ula30c]. Massproblem [Ula30a]. Masstheorie [Ula30d]. masters [Ano82, Ano93]. Matching [Dub76]. Matematika [KU71]. matematyka [Ula78c]. Math [Wil76, Ula67b]. Mathematical [AA85, AA08, GM61, MU87b, RB12, SUW64, SU67b, Ula60a, Ula61b, Ula62b, EU48b, Jun01, TWHM81, UP53, Ula56d, Ula62d, Ula78d, Ula80c, Ula81c, Ula85, Ula86, BU90, Ula13]. mathematically [Ula66b, Ula67a]. Mathematician [Ula87d, Ula76a, Ula91a, Met76, Wil76, Bir77]. Mathematicians [Ano15, Ula76b]. Mathematics [Ano87e, Bec62, Bec13, CR09, GM61, KU68, KU79, KU92, Kem70, MOR76a, Pat87, PZHC09, Ula80d, URR86a, URR86e, FF91, Gar01a, Kur80, Mau81, MOR76b, Moo82, RR82, Ula64c, Ula69b, Ula72c, Ula81a, URR86d, TWHM81, Mau15, Asp88, Ula47b, May69]. Mathématiques [KU73]. Matousek [Bár04]. matrices [EU50d]. matrix [SZ10]. matters [Rei72]. May [Ano05, DKU85, Fla84]. Meaning [Rot86a, Rot85, Rot89]. means [EU55]. Measurable [Myc87]. Measure [Gra87, OU41, BU42, LU34, Myc77, OU38a, OU39, Ula30a, Ula30b, Ula30d, Ula30c, US31, Ula32, Ula38b, Ula39b, Ula41a, Ula43a, Ula97]. Measure-preserving [OU41]. measures [Ula41a]. mechanics [Rot86b]. mechanism [Ula82e]. Meeting [Ano62, UdvB^{+}64]. Meetings [BHMU85, Hur85]. Memo [MU87a]. Memoir [Haw87]. Memorable [MU87a]. memorial [Rei72, Rot84]. Memoriam [Rot86a]. memorium [Rot89]. Memory [AMS19, Fla85, RR82]. Mengen [Ula31, Ula33b]. Mengenlehre [Ula30d]. mesure [LU34, US31]. Method [Eck87, Met87, Ben16, EU55, FS10, Fro99a, Fro99b, MU49b, Ula51, Ula61c]. Methods [Bár04, Bud79, KW64, Ula56b, RUvN47, Ula48, Ula56a]. Metric [BU79, BSSU74, Ula38b, Ula43c]. metrical [OU41]. metrically [OU38b]. Metrics [BSSU72]. Mexico [Ano05, DKU85]. migration [FB69]. Milestones [Gon96a, Gon96b]. Mind [FE85, Ney74]. Minimal [CEG^{+}79]. Minkowski [Ula43c]. Minneapolis [Ula71b]. Minnesota [Ula71b]. Mirrors [TU51]. Mitchell [Ano93]. mixing [Ula82f]. Mixture [EU50c, EU50b]. Model [AMS19, Ula67c]. modeling [US71, US86]. models [Ula57a]. Modern [Bec62, Bec13, Mac92, McM05, Ula64c]. Molecular [BSSU74, Goa87]. Monte [JML13, Ben16, BHMU85, Dys12a, Eck87, HPR14, Hur85, KW64, MU49b, Ula56b, Met87, Ula48, Ula51, Ula56a, Ula61b, Ula62d, Ula13]. most [Har10]. Mostowski [Ula47b]. motions [MU49a]. Much [Mac92]. multiplication [Ula61c]. Multiplicative [EU48a, EU49a, EU49b, EU49c, UE48a, UE48b, UE48c, UH44]. My [Gam70b, Gam70a]. Mysteries [LZ12]. mysterious} [Dau08].

n [SU35a]. Name [Cip00]. Narodziny [Ula78b]. National [DKU85]. Natural [MOR76a, SU33b, SU37, US33, Ula78d, Ula81a]. naturels [US33]. natürlichen [SU33b, SU37]. navigating [Ula58e]. networks [SZ10]. neue [UB31]. Neumann [Bar86, Ber10, Eck87, JML13, Mac92, Réd05, RW12, Ula58b, Ula58d, UKTS68, UKTS69, Ula80d, Ula82a, Ula82b, URR86b, URR86c, URR86e]. neutron [RUvN47]. never [Kno90]. nombre [AU35, SU35b]. nombres [US33]. Non [SU63, SU64, Ula63c]. Non-linear [SU63, SU64, Ula63c]. Nonlinear [Ano05, Cam87, DKU85, Mau87, SU73, FPU55, PUF65]. normed [MU31]. normés [MU32]. normierten [MU31]. Note [BU68, Hur85]. Notes [SUW64, SU67b]. notion [BSU71, MU69, Ula81c, Ula86]. notions [Ras10]. November [TWHM81, UdvB^{+}64]. Nuclear [EU61, HPR14, Mac92, EU55, UdvB^{+}64, Ula63a]. Number [GM61, Ste87, AU35, Gar01a, Rot86b, SU33b, SU35b, SU37]. numbers [BMR74, EU50d, LU33a, US33]. Numerical [Har87, PU59, Ula68b, DPR05, PZHC09, US71, US86]. NY [UdvB^{+}64].

objects [SU67a]. Observation [SU67b]. October [Ano62, TWHM81]. odyssey [Yor87]. old [Ano15, Ula74b]. one [Har10]. open [Ula62e]. operation [UB31, Ula38d, Ula41b, Ula43d, Ula44, UB31]. operations [Ula80a]. Oppenheimer [McM05, PC06]. Optimal [Dub76]. ordered [EU44b, EU45a]. organization [Ula57c]. origins [Dys12b]. Orion [Dys02, Ula65a]. other [Fro99b, Gar01a, Rei72]. our [Hua76]. Outline [Hol62]. Oxford [Moo82].

P [Str98]. P34 [BU47]. P35 [BU47]. Packing [Gar01b, LZ12]. pair [Ula80a]. pairing [MU69]. pairwise [CEG^{+}79]. panel [UdvB^{+}64]. Papers [Ano62, TWHM81]. Paradigms [Cam87]. paradoxes [Gar01a]. parallel [BU77a]. Part [MSU59, UE48a, UE48b, UE48c]. Paso [TWHM81]. Pasta [AMS19, Ano05, Ano15, BLP09, BP11, BCP13, BPP18, BI04, BI05, DPR05, OVPL15, Str98, Wei97, Dau08, PZHC09]. Patterns [BU79, SU60, Ula62b, HU60, SU67a, Ula65b]. peace [Yor87]. pedagogical [DPR05]. People [AA85, Asp88, AA08, URR86d]. Pergamon [Moo82]. period [KW64]. permutation [SU33b, SU37, US33]. permutations [US33]. Permutationsgruppe [SU33b, SU37]. Personal [URR86a, For15]. perspectives [DPR05, KU73]. perspektivy [KU71]. Philosophical [Ula68c]. photographs [Moo82]. Physical [Ula64a, Ula85]. Physicist [Gol99, Ula80c, Yor87]. physicists [Har06]. Physics [Ano87f, Hua76, Pat87, Ula61b, Ula76b, FF91, RR82, UP53, Ula57a, Ula57b, Ula62d, Ula13]. physics-astrophysics [RR82]. Pierre [Ula75]. Pioneered [Mac92]. Play [Haw87]. pleasing [Ney74]. points [MU49a, Ula30e, Ula47a]. Pointwise [Bar86]. Poland [Kar77]. Polish [Moo82, Ano82, Ano93, Kur80, Ula83, Zie77]. populations [EU69]. positive [EU50d]. possibilities [EU66, Ula57c]. Possibility [UW64, Ula58e]. possible [Ula78d, Ula85]. Potpourri [Ste87]. Poznan [Kar77]. Poznan-Kornik [Kar77]. pp [Bár04, Bir77, Moo82]. Practicalities [Cam87]. Practice [BU77b]. Preface [Moo82]. Preliminary [BU75, Ula42, EU44a, HU41, OU38b, Ula38a, Ula39a, Ula41b, Ula43c, Ula43d, Ula44]. preserving [OU41]. Primes [SUW64, SU67b]. Princeton [Ano62]. probabilistic [EU71]. probabiliteés [LU34]. probabilities [LU34, Ula57b]. Probability [Mau87, Rot86b, Ula49, Gar01a, LU33b]. Problem [Ano05, Ano15, BLP09, BP11, BCP13, BPP18, Dub76, Pat87, Wei97, BI04, BI05, BU42, EU48b, RW12, Ula30a, Ula43c, Ula62e, EU50c, EU50b, Ula33a, Str98]. Probleme [Ula38c, Ula33a]. Problemes [BU47]. Problems [Mau15, PU59, Ula49, Ula60a, Ula61b, Ula62b, Ula64c, CU44, EU44a, FPU55, Gar01a, HU60, MU87b, PUF65, SU55, Ula48, UP53, Ula56d, Ula61c, Ula62d, Ula63b, Ula66b, Ula67a, US71, Ula72a, US86, Ula13, GM61]. proceedings [BC82, Bud79, CFK71, Kar77, MOR76a, MOR76b, Ano51, DKU85]. process [MU69, UW64]. processes [BSU69, UH44, UvN47, Ula52, Ula68a, US71, US86]. product [Ula32, UK32, Ula38d, Ula44, Ula71a]. production [Ula80a]. products [BU31, Ula41b, Ula43d]. produit [UK32]. Produktraumen [Ula32]. Profiles [AA85, AA08]. progress [BI04, BI05]. project [Ula65a, Ben16, Dys02]. projectiles [EU55]. Projective [BU75, BU78, EU45b, EU46]. Propagation [DKU85]. propelled [EU61]. Properties [SUW64, UK32, Ula61a, Ula63c]. property [AMU35, MU53, MUxx, US31]. proposition [Ula31]. Propositions [Ula47b]. propriété [AMU35, US31]. propriétés [UK32]. propulsion [EU55]. Prospect [May69]. prospects [KU68, Ula72b, Kem70]. Proud [Zie77]. Przygody [Ula78c]. Publications [Ano87a]. Publishers [Moo82]. Pure [GM61]. Pursuing [Str98, Wei97]. puzzles [Gar01a]. PWN [Moo82].

Quadratic [MSU59]. quantities [Ula31, Ula33b]. quasi [Ula47a]. quasi-fixed [Ula47a]. Quelques [UK32]. Questions [Zor24].

R1 [BU47]. Race [Gon96b, McM05]. Radiation [TU51]. Radisson [Ano05]. Rand [Ula71b]. Random [EU50e, Ula52, vNU45, BSU69, Fro99a, Ula80a]. randomly [MU49a]. randomness [MU53, MUxx]. Rassias [Jun01]. rate [Ula66b, Ula67a]. rates [US71, US86]. Raumen [SU34b, Ula30c]. reaction [SU59]. reactions [EU50a]. Real [Ula59a, Ula59b, EU50d, Goo04, Ula38f]. Really [HM90]. recognition [Ula82e]. recreational [Gar01a]. Recursive [Ula65b]. recursively [SU67a]. Reflections [Ano82, Ano93, Arr89, Gra87, Rid89, Ula87c, CES89, Kur80, Moo82]. regard [RW12]. Regular [LZ12]. related [UK33]. relation [BU42]. Relational [BU77b, BU77a]. Relations [BU75, Ula63a, BU76, BU78, Ula63a]. Relativity [EU66]. Remark [Ula29b, SU35a, SU34b]. Remarks [BU77b, EU71]. Remembrances [Moo82, Kur80]. reminiscence [Ula83]. Reminiscences [Ula69a, Ula69c, URR86a, Ula94]. Rencontres [CU^{+}66]. Reply [Rot90]. Report [BU75, Wel12, EU44a, HU41, OU38b, Ula38a, Ula39a, Ula41b, Ula42, Ula43c, Ula43d, Ula44]. reported [PZHC09]. reports [BU90]. representations [Ula42]. Representatives [Ula58f]. Reproducing [vNB66]. Research [Ula60b, Ula62c, Ula60b]. Retrospect [Kem70, May69, KU68]. rétrospective [KU73]. Retrospektiva [KU71]. Review [Asp88, Bár04, Bir77, GM61, May69, Met76, Moo82, Str98, Ula47b, Ula59a, Ula59b]. Reviews [Kem70]. Reynolds [Asp88]. Richard [Zor24]. Rietz [Bud79]. Ritzen [Ula75]. Robert [McM05]. robot [CU^{+}66]. rocket [EU61]. role [Ula78d]. Roman [Ula59a, Ula59b]. Rosa [Ula87e]. Rota [Asp88, Ula87a]. Route [OVPL15]. ruin [McM05]. rules [Ula74b]. Rzeczywiste [Ula59b, Ula59a].

S [Bir77, Met76, MOR76a, Wil76]. S. [Dub76, GM61, MOR76b, Rot84, BU90, Ula94]. Sandwiches [Hil18]. Santa [Ano05]. Satze [Ula31]. Sausage [Zor24]. Scale [Ano51]. Scales [BP11]. Scaling [BLP09]. schemata [BU77a, Ula81c, Ula86]. Science [Cam87, URR86d, RR82, Har06, Asp88]. sciences [Ula78d, Ula81a, MOR76a]. Scientific [BHMU85, Hur85, Mac92, Moo82, Ula62c, Ula68c]. scientists [Har10]. Scottish [Mau81, Mau15, Mau15, Ula57d, Ula69a, Ula69c, Ula77, Ula82c]. Scribners [Bir77]. Second [Ano51, Bec62, Bec13, Ula31]. Secret [HM90, Kno90, WG17]. secrets [HM97]. Segments [Dub76]. selbst [SU35a]. Selected [Mau15, RR82, BMR74, Réd05]. Self [vNB66]. Self-Reproducing [vNB66]. semi [AU35]. semi-simple [AU35]. semisimple [AU35]. Senate [Ula63a, Ula63a]. separable [Ula38b]. September [Ano51, Bud79, Kar77, Ula71b]. Sequence [BSSU74, Goa87, Wei06]. sequences [GLMU56, SU33b, SU37, US33]. Series [Bec62, Bec13]. service [Rot84]. Set [Car03, Ula38d, BU68, EU44a, OU38a, Ula30d, Ula97]. Set-theoretical [Ula38d]. Sets [Ula64a, BU42, BMR74, EU68, MU67, Ula29a, UK33, Ula41a, Ula41b, Ula43d, Ula44]. several [UE48a, UE48b, UE48c]. shock [Han92]. sich [SU35a]. sieves [GLMU56]. Sikorski [Ula59a, Ula59b]. Similarity [Gra87]. simple [AU35]. Simulation [Wei97, Str98, Ula67c]. Simulations [HPR14]. Size [BP11]. Slicing [Hil18]. Smoluchowski [Ula57b]. Society [UdvB^{+}64]. Solar [Hil18]. Solution [Dub76]. Solve [Ano15]. Solvers [JML13]. solving [Ula61c]. Some [BU77b, EU44a, EU71, PU59, SUW64, Ula56d, Ula62b, Ula63c, Ula63b, Ula64a, US71, Ula72a, Ula72b, US86, BSU69, CU44, EU66, HU60, Ula57c, Ula61a, Ula68c, Ula81c, Ula85, Ula86, UK32]. son [LU34]. Soviet [WG17].

Space [Ula58f, HU42, HU47, Ula38a, Ula38b, Ula58e, UdvB^{+}64, Ula73]. space-time [Ula73]. spaces [BU31, LU34, MU31, SU34b, Ula30c, Ula32, Ula47a]. spaceship [Dys02]. special [RW12, EU66]. specifically [Ula78d]. Speculations [Ula82e, Ula85]. speed [UP53]. Sperry [Ula71b]. Sphären [SU34a]. Sphere [vNU47, SU34a]. spheres [Gar01b, SU33a, SU33a]. Spirit [Haw87]. spoken [Rot84]. sponsored [UdvB^{+}64]. Springer [Bár04]. Stability [Ula62a, Jun01, Ras10, UH54]. Stage [Car03]. Stan [Ano82, Ano93, Bar81, CR09, Eck87, Fei85, Haw87, Mar87, Mys90, Rot86a, Rot87, Rot90, Ula75, Ula87b, Ula87d, Wat06]. standard [FS10]. Stanislaw [Asp88, Kem70, May69, Moo82, Ano87b, Ano87a, Ano87c, Arr89, Bar85, Bar08, BMR74, BSW85, CES89, FE85, Fla84, Fla85, Hoa02, Myc90, PZHC09, Rid89, Rot89, Sul84]. star [Ula68b]. stars [UW64]. Statement [Ula58f, Ula60b]. States [Hil18, Ula63a, Ula71b]. static [Ula65b]. Statistical [RUvN47, Ula48, Rot86b, Ula61c, Ula61a]. Stefan [Ula46]. stellar [Ula66a]. stetige [Ula34]. stetigen [SU34b]. stochastic [UvN47]. story [Dys02]. Strange [Ste87]. Strangelove [Goo04]. structures [BU77a, EU48b]. studied [Ula72a]. Studies [Ano05, FPU55, PUF65, BSU69, RR82, SU63, SU64, UP53, Ula68b]. Study [SU55]. Sub [Ula87e]. Subcommittee [Ula60b]. subgraphs [CEG^{+}79]. subsets [Ula41a]. such [EU61]. suite [US33]. Surface [vNU47, Ula43c]. Surveys [MOR76b]. symmetric [BU31]. Symposium [Ano51, MOR76a, Hua76, BC82]. system [OVPL15]. Systems [DKU85, Hil18, Hol62, Mau87, Ula47b, BC82, CU44, EU48a, EU49a, EU49b, EU49c, MU49a, RW12, SU59, UE48a, UE48b, UE48c, Ula56c, Ula58e, Ula61a, Ula66a, Ula68b]. Szkockiej [Ula69c, Ula78b].

T [EU50b]. tactical [Ula56b, Ula56a]. talking [Yor87]. Technology [Hua76]. Teller [Goo04, Har10, Wel12]. Test [Ula63a]. Testimony [Ula71b]. Tetrahedra [LZ12]. Texas [TWHM81]. their [EU48b]. Theme [Gra87]. Theorem [Bár04, EU71, Ula29b]. Theorems [vNU45]. Theoretical [BSW85, Ula38d]. Theorie [Ula30e, LU34]. Theories [Ula64a, Ney74, Ula85]. Theorist [Sul84]. Theory [Ano62, BU77b, BSW85, Hol62, Mac92, Ste87, Ula38f, Ula41b, Ula43d, UH44, Ula44, vNB66, BU77a, Bud79, CFK71, EU44a, Gar01a, Kar77, LU34, Rot86b, Ula30d, Ula30e, Ula39b, Ula42, Ula57b, Ula97, EU66]. thermalization [OVPL15]. Thermodynamic [BP11]. Thermonuclear [Gon96a, Gon96b, Ula66d, EU50a]. third [TWHM81]. Thomas [Str98]. Time [BP11, Ula73]. Time-Scales [BP11]. Top [Cip00]. topics [Gar01a]. Topological [Bár04, BU31, SU34b, SU34a, SU35a, SU35b, UB31, UK32, Ula42]. topologique [SU35b]. topologiques [UK32]. topologische [SU34a, UB31]. Topologischen [SU34b, SU35a]. topology [Gar01a]. Touch [Ano87g]. tower [Wil76]. Tracts [GM61]. transformation [OU39, SU63, SU64].

Transformations [SU73, Ula82f, MU31, MU32, MSU59, OU38b, SU33a, SU34b, SU34a, SU35a, UK33, Ula33a, Ula38a, Ula38e, Ula39a, Ula47a, Ula52, Ula63c]. transitive [OU38b]. transitivity [OU41]. transmutations [Ula80a]. Transport [DKU85]. treasury [FF91]. Treaty [Ula63a]. Tree [Asp88, URR86d]. Trees [BSSU74]. trends [CFK71]. Trialogue [Ula87e]. Tritium [EU50c]. true [Dys02]. Turbulence [BHJ^{+}87]. Turing [Dys12b]. twentieth [Har06, Har10, MHR80]. Two [SU60, CEG^{+}79, Ula38f]. type [Ula48].

U.S. [Ula58f, Ula63a]. Ulam [Ano87b, Ano05, Asp88, Bár04, Bir77, Fla84, GM61, Gle02, Jun01, MOR76a, Moo82, OVPL15, Str98, AMS19, Ano82, Ano87a, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano93, Ano15, Arr89, Bar86, Bar81, Bar85, Bar08, BLP09, BP11, BCP13, BPP18, BI04, BI05, BMR74, BSW85, Car03, CR09, CES89, DPR05, Dau08, Dub76, Dys12c, Eck87, Fei85, FE85, Fla85, FS10, Fro99a, Fro99b, Goa87, Gra87, Han92, Haw87, Hoa02, JML13, Kem70, Mar87, May69, Met76, MOR76b, Myc77, Myc87, Myc90, Mys90, PZHC09, Ras10, Rid89, Rot84, Rot86a, Rot87, Rot89, Rot90, SZ10, Sul84, Ula75, Ula87b, Ula87d, BU90, Ula94, Wat06, Wei97, Wei06, Wel12, Wil76]. Ulamian [Ste87]. Undecidable [Ula47b]. Underlying [BCP13, BPP18]. Understand [Ula87c]. unendliche [MU30]. Uniqueness [Ula30b]. United [Ula63a, Ula71b]. universe [Dys12b]. universes [BMR74]. University [Ano51, Ano62, BC82, TWHM81, Ula38f]. unknowns [EU68]. Unlikely [Ber10]. USA [Ano05]. use [Ula57c]. uses [Ula58c]. Using [Bár04, Fro99b, JML13].

Valued [BU79]. Variables [LU34, Ula38f, UE48a, UE48b, UE48c]. vector [MU31]. vectoriels [MU32]. vectors [CU44]. vehicle [EU61]. vehicles [Ula58e]. Vektorraumen [MU31]. versus [Ula71b]. very [UW64]. Victor [Hua76]. viii [Moo82]. Visual [SUW64, BU68, MU67]. Vita [Ula87d]. vollstandig [Ula30c]. Volume [Ula59b, RR82, Rei72, Rot86b]. vs [BLP09].

Wahrscheinlichkeitsrechnung [LU33b]. walk [EU50e, Ula80a]. Warmth [FE85]. Warsaw [Moo82]. Was [Mys90, Rot90, KAS96, Kno90]. waves [Han92]. way [KAS96]. weapons [Yor87]. Weissert [Str98]. Weisskopf [Hua76]. Wendisch-Rietz [Bud79]. Who [HM90, Mac92, Har06]. winter [UdvB^{+}64]. Words [Rot84]. Work [Har87, PU59]. works [BMR74]. World [Gam70b, Gam70a, FF91, Hua76]. Wspomnienia [Ula69c, Ula83].

year [Ano15]. Years [Ano05, BI04, BI05, PZHC09]. York [Bir77, CFK71, GM61, UdvB^{+}64].

Zahlen [LU33a]. Zahlenfolge [SU33b, SU37]. Zerlegungen [Ula33b]. zero [OU38a, Ula41a]. zum [Ula30a, Ula32]. Zur [Ula30d, Ula30e].

References

Albers:1985:MPP

[AA85] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical People: Profiles and Interviews. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1985. ISBN 0-8176-3191-7, 3-7643-3191-7. xvi + 2 + 372 pp. LCCN QA28 .M37 1985.

Albers:2008:MPP

[AA08] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical people: profiles and interviews. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2008. ISBN 1-56881-340-6. xxvi + 386 pp. LCCN QA28 .M37 2008.

Amati:2019:MEF

[AMS19] Graziano Amati, Hugues Meyer, and Tanja Schilling. Memory effects in the Fermi–Pasta–Ulam model. Journal of Statistical Physics, 174(1):219–257, January 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Auerbach:1931:CEE

[AMU31] H. Auerbach, S. Mazur, and Stanisław M. Ulam. Über eine charakteristische Eigenschaft des Ellipsoides. (German) [On an eigenvalue of ellipsoids]. Annales de la Société Polonaise de Mathématiques, 10(??):128–??, ???? 1931.

Auerbach:1935:PCE

[AMU35] H. Auerbach, S. Mazur, and S. Ulam. Sur une propriété caractéristique de l’ellipsoïde. (French) [On a characteristic property of the ellipsoid]. Monatshefte für Mathematik und Physik, 42(1):45– 48, 1935. CODEN MMPHA8. ISSN 0026-9255 (print), 1436-5081 (electronic).

Anonymous:1951:PSS

[Ano51] Anonymous, editor. Proceedings of a Second Symposium on Large-Scale Digital Calculating Machinery, Harvard University Computation Laboratory, 13–16 September 1949, volume 26 of Annals of the Computation Laboratory of Harvard University. Harvard University Press, Cambridge, MA, USA, 1951. LCCN QA75 .S9 1949.

Anonymous:1962:RAG

[Ano62] Anonymous, editor. Recent Advances in Game Theory: Papers Delivered at a Meeting of the Princeton University Conference, October 4–6, 1961, volume 29 of Princeton Conference Series. Princeton University Press, Princeton, NJ, USA, 1962. LCCN QA269.P75.

Anonymous:1982:RPM

[Ano82] Anonymous. Reflections of the Polish masters: An interview with Stan Ulam and Mark Kac. Los Alamos Science, 3(3):54–65, 1982. CODEN LASCDI. ISSN 0273-7116.

Anonymous:1987:PSM

[Ano87a] Anonymous. Publications of Stanisław M. Ulam. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):313–318, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://library.lanl.

gov/cgi-bin/getfile?15-26.pdf.

Anonymous:1987:SU

[Ano87b] Anonymous. Stanisław Ulam, 1909–1984, volume 15 of Los Alamos science. Los Alamos National Laboratory, Los Alamos, NM, USA, 1987. 318 pp. LCCN Q1 .L67; Q11 .L5.

Anonymous:1987:SUC

[Ano87c] Anonymous. Stanisław Ulam: Cover and contents. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):c1, c2, 1–3, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://library.

lanl.gov/cgi-bin/getfile?15-00.pdf.

Anonymous:1987:ULB

[Ano87d] Anonymous. The Ulam legacy: Biology. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):280–282, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-19.pdf.

Anonymous:1987:ULM

[Ano87e] Anonymous. The Ulam legacy: Mathematics. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):34–38, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-05.pdf.

Anonymous:1987:ULP

[Ano87f] Anonymous. The Ulam legacy: Physics. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):122–124, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-11.pdf.

Anonymous:1987:UT

[Ano87g] Anonymous. The Ulam touch. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):292–293, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-22.pdf.

Anonymous:1993:RPM

[Ano93] Anonymous. Reflections of the Polish masters: An interview with Stan Ulam and Mark Kac by Mitchell Feigenbaum. Annales Societatis Mathematicae Polonae. Seria II. Wiadomości Matematyczne, 30(1):93–114, 1993. ISSN 0373-8302.

Anonymous:2005:CNS

[Ano05] Anonymous, editor. The Center for Nonlinear Studies 25th Annual International Conference: 50 Years of the Fermi–Pasta–Ulam Problem: Legacy, Impact, and Beyond: May 16–20, 2005, Radisson Santa Fe Hotel, Santa Fe, New Mexico USA. ????, ????, 2005. ISBN

???? LCCN ???? URL http://cnls.lanl.gov/Conferences/
annual25/agenda.htm.

Anonymous:2015:MSY

[Ano15] Anonymous. Mathematicians solve 60-year-old Fermi–Pasta– Ulam problem. Scientific Computing, ??(??):??, March 24, 2015. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156

(electronic). URL http://www.scientificcomputing.com/news/
2015/03/mathematicians-solve-60-year-old-fermi-pasta-ulam-

problem. See [OVPL15] for the research paper.

Arrowsmith:1989:CCR

[Arr89] D. K. Arrowsmith. From cardinals to chaos — reflections on the life and legacy of Stanisław Ulam. Nature, 342(6246):137, November 9, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Aspray:1988:BRB

[Asp88] William Aspray. Book review: Science, Computers, and People: From the Tree of Mathematics by Stanislaw Ulam; Mark C. Reynolds; Gian-Carlo Rota. Isis, 79(4):702–703, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234776.

Auerbach:1935:NGG

[AU35] H. Auerbach and S. Ulam. Sur le nombre de générateurs d’un groupe semi-simple. (French) [On the number of generators of a semisimple group]. Comptes rendus de l’Académie des sciences, Paris, 201:117–119, 1935.

Barcellos:1981:ISU

[Bar81] Anthony Barcellos. An interview with Stan Ulam. Two-Year College Mathematics Journal, 12(3):182–189, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/00494925.1981.11972521.

Barcellos:1985:SMU

[Bar85] Anthony Barcellos. Stanisław M. Ulam. In Albers and Alexanderson [AA85], pages 367–374. ISBN 0-8176-3191-7, 3-7643-3191-7. LCCN QA28 .M37 1985.

Barbour:1986:PCD

[Bar86] A. D. Barbour. Pointwise convergence of densities under iteration of Ulam and von Neumann’s map. Journal of Physics A (Mathematical and General), 19(18):3921, 1986. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http://stacks.

iop.org/0305-4470/19/i=18/a=037.

Barany:2004:BRU

[Bár04] Imre Bárány. Book review: “Using the Borsuk–Ulam Theorem: Lectures on Topological Methods in Combinatorics and Geometry,” by J Matoušek, Springer, 2003, 196 pp., £30.50, $49.95, EUR49.95. Combinatorics, Probability and Computing, 13(2):281– 282, March 2004. CODEN CPCOFG. ISSN 0963-5483 (print),

1469-2163 (electronic). URL http://journals.cambridge.org/
action/displayIssue?jid=CPC&volumeId=13&issueId=02.

Barcellos:2008:BM

[Bar08] Anthony Barcellos. Stanisław M. Ulam. In Albers and Alexanderson [AA08], pages 367–374. ISBN 1-56881-340-6. LCCN QA28 .M37 2008.

Bednarek:1982:DSI

[BC82] A. R. Bednarek and Lamberto Cesari, editors. Dynamical systems II: proceedings of a University of Florida international symposium. Academic Press, New York, NY, USA, 1982. ISBN 0-12-084720-5, 0-12-083750-1. LCCN QA614.8 .D93 1982.

Benettin:2013:FPU

[BCP13] G. Benettin, H. Christodoulidi, and A. Ponno. The Fermi–Pasta– Ulam problem and its underlying integrable dynamics. Journal of Statistical Physics, 152(2):195–212, July 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-013-0760-6.

Beckenbach:1962:MME

[Bec62] Edwin Ford Beckenbach, editor. Modern Mathematics for the Engineer: Second Series. McGraw-Hill, New York, NY, USA, 1962. xviii + 456 pp. LCCN QA401 .B4. With an introduction by Magnus R. Hestenes.

Beckenbach:2013:MMEb

[Bec13] Edwin Ford Beckenbach, editor. Modern Mathematics for the Engineer: Second Series. Dover Publications, Inc., New York, NY, USA, 2013. ISBN 0-486-31612-2 (e-book), 0-486-49747-X (paperback), 1-306-32668-0 (e-book). xviii + 456 pp. LCCN QA401 .B39 2013. With an introduction by Magnus R. Hestenes. Reprint of [Bec62].

Benov:2016:MPF

[Ben16] Dobriyan M. Benov. The Manhattan Project, the first electronic computer and the Monte Carlo method. Monte Carlo Methods and Applications, 22(1):73–??, March 2016. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL

https://www.degruyter.com/view/j/mcma.2016.22.issue-1/
mcma-2016-0102/mcma-2016-0102.xml.

Bernstein:2010:JNK

[Ber10] Jeremy Bernstein. John von Neumann and Klaus Fuchs: an unlikely collaboration. Physics in Perspective (PIP), 12(1):36–50, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0001-1;   http://www.springerlink.com/
content/y54804274h88v546/.

Besnard:1987:IT

[BHJ^{+}87] Didier Besnard, Francis H. Harlow, Norman L. Johnson, Rick Rauenzahn, and Jonathan Wolfe. Instabilities and turbulence. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):145– 174, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://

library.lanl.gov/cgi-bin/getfile?15-15.pdf.

Bowes:1991:LMF

[BHM91] Stanley A. Bowes, Daniel Hirsch, and William G. Mathews. Letters: More fallout from H-bomb controversy. Bulletin of the Atomic Scientists, 47(9):45–46, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90] and comment [Han92].

Burks:1985:CEM

[BHMU85] Arthur W. Burks, Alston S. Householder, N. Metropolis, and S. M. Ulam. Comments on early Monte Carlo computations and scientific meetings. Annals of the History of Computing, 7(2):147– 148, April/June 1985. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1985/pdf/a2141.pdf;
 http://www.computer.org/annals/an1985/a2141abs.htm.

Berman:2004:FPU

[BI04] G. P. Berman and F. M. Izrailev. The Fermi–Pasta–Ulam problem: 50 years of progress. arxiv.org, ??(??):1–48, November 29, 2004.

URL http://arxiv.org/abs/nlin/0411062v3.

Berman:2005:FPU

[BI05] G. P. Berman and F. M. Izrailev. The Fermi–Pasta–Ulam problem: Fifty years of progress. Chaos (Woodbury, NY), 15(1):015104, 2005. CODEN CHAOEH. ISSN 1054-1500. URL http://chaos.aip.

org/resource/1/chaoeh/v15/i1/p015104_s1.

Birkhoff:1977:BRB

[Bir77] Garrett Birkhoff. Book review: Adventures of a mathematician: By S. M. Ulam. New York (Scribners). 1976. 317 pp. $17.50. Historia Mathematica, 4(2):233–236, May 1977. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL http://www.

sciencedirect.com/science/article/pii/0315086077901318.

Benettin:2009:FPU

[BLP09] G. Benettin, R. Livi, and A. Ponno. The Fermi–Pasta–Ulam problem: Scaling laws vs. initial conditions. Journal of Statistical Physics, 135(5–6):873–893, June 2009. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-008-9660-6.

Beyer:1974:SUS

[BMR74] W. A. Beyer, J. Mycielski, and Gian-Carlo Rota. Stanisław Ulam: sets, numbers, and universes: selected works, volume 9 of Mathematicians of our time. MIT Press, Cambridge, MA, USA, 1974. ISBN 0-262-02108-0. xxiii + 709 pp. LCCN QA3 .U4; QA3 .U41; QA36 .U36s.

Benettin:2011:TSE

[BP11] G. Benettin and A. Ponno. Time-scales to equipartition in the Fermi–Pasta–Ulam problem: Finite-size effects and thermodynamic limit. Journal of Statistical Physics, 144(4):793–812, August 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/10.
1007/s10955-011-0277-9.

Benettin:2018:FPU

[BPP18] G. Benettin, S. Pasquali, and A. Ponno. The Fermi–Pasta–Ulam problem and its underlying integrable dynamics: An approach through Lyapunov exponents. Journal of Statistical Physics, 171 (4):521–542, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Beyer:1972:MBI

[BSSU72] W. A. Beyer, M. L. Stein, Temple Smith, and Stanisław M. Ulam. Metrics in biology, an introduction. Technical Report LA-4973, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1972.

Beyer:1974:MSM

[BSSU74] W. A. Beyer, M. L. Stein, T. F. Smith, and S. M. Ulam. A molecular sequence metric and evolutionary trees. Mathematical Biosciences, 19(1–2):9–25, 1974. CODEN MABIAR. ISSN 0025-5564 (print), 1879-3134 (electronic).

Beyer:1969:CSS

[BSU69] W. A. Beyer, R. G. Schrandt, and Stanisław M. Ulam. Computer studies of some history-dependent random processes. Technical Report LA-4246, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1969.

Beyer:1971:NC

[BSU71] W. A. Beyer, M. L. Stein, and Stanisław M. Ulam. The notion of complexity. Technical Report LA-4822, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1971.

Beyer:1985:USM

[BSW85] W. A. Beyer, P. H. Sellers, and M. S. Waterman. Stanisław M. Ulam contributions to theoretical theory. Letters in Mathematical Physics, 10(2–3):231–242, ???? 1985. CODEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Borsuk:1931:SPT

[BU31] Karol Borsuk and Stanisław M. Ulam. On symmetric products of topological spaces. Bulletin of the American Mathematical Society, 37(12):875–882, ???? 1931. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.

org/euclid.bams/1183495161.

Bernstein:1942:PCA

[BU42] D. L. Bernstein and Stanisław M. Ulam. On the problem of completely additive measure in classes of sets with a general equivalence relation. Bulletin of the American Mathematical Society, 48 (??):361–362, ???? 1942. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Banach:1947:PPP

[BU47] Stefan Banach and Stanisław M. Ulam. Problemes P34, P35, P35,R1. Colloquium Mathematicum, 1(??):152–153, ???? 1947. CODEN CQMAAQ. ISSN 0010-1354 (print), 1730-6302 (electronic).

Beyer:1968:NVH

[BU68] W. A. Beyer and S. M. Ulam. Note on the visual hull of a set. Journal of Combinatorial Theory, 4:240–245, 1968. CODEN JCTHAR. ISSN 0021-9800 (print), 1878-1756 (electronic).

Bednarek:1975:PAC

[BU75] A. R. Bednarek and S. M. Ulam. Projective algebras and calculus of relations — preliminary report. Notices of the American Mathematical Society, 22(7):A729, ???? 1975. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bednarek:1976:GAR

[BU76] A. R. Bednarek and S. M. Ulam. Generators for algebras of relations. Bulletin of the American Mathematical Society, 82(5):781– 782, 1976. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-

881X (electronic). URL http://projecteuclid.org/euclid.
bams/1183538242.

Bednarek:1977:TRS

[BU77a] A. R. Bednarek and Stanisław M. Ulam. On the theory of relational structures and schemata for parallel computation. Technical Report LAMS-6734, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1977.

Bednarek:1977:SRR

[BU77b] Alexander R. Bednarek and S. M. Ulam. Some remarks on relational composition in computational theory and practice. In Karpinski [Kar77], pages 22–32. CODEN LNCSD9. ISBN 0-387-08442-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1,L43 no.56. URL http://link.springer-ny.

com/link/service/series/0558/tocs/t0056.htm;   http://
www.springerlink.com/openurl.asp?genre=issue&issn=0302-
9743&volume=56.

Bednarek:1978:PAC

[BU78] A. R. Bednarek and S. M. Ulam. Projective algebra and the calculus of relations. Journal of Symbolic Logic, 43(1):56–64, 1978. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL

http://projecteuclid.org/euclid.jsl/1183740106.

Bednarek:1979:IVM

[BU79] A. R. Bednarek and S. M. Ulam. An integer-valued metric for patterns. In Budach [Bud79], pages 52–57. ISBN ???? LCCN QA267 .C594 1979.

Ulam:1990:ABA

[BU90] A. R. Bednarek and Françoise Ulam, editors. Analogies between analogies: the mathematical reports of S. M. Ulam and his Los Alamos collaborators, volume 10 of Los Alamos Series in Basic and Applied Sciences. University of California Press, Berkeley, CA, USA, 1990. ISBN 0-520-05290-0. xviii + 565 pp. LCCN QA274 .A53 1990; Internet. The mathematical reports of S. M. Ulam and his Los Alamos collaborators, Edited and with a foreword by A. R. Bednarek and Françoise Ulam, With a bibliography of Ulam by Barbara Hendry.

Budach:1979:FCT

[Bud79] L. (Lothar) Budach, editor. Fundamentals of computation theory: FCT ’79: proceedings of the Conference on Algebraic, Arithmetic, and Categorial Methods in Computation Theory held in Berlin/ Wendisch-Rietz (GDR), September 17–21, 1979, volume 2 of Mathematical research. Akademie-Verlag, Berlin, Germany, 1979. ISBN ???? LCCN QA267 .C594 1979.

Burks:1970:ECA

[Bur70] Arthur W. Burks. Essays on Cellular Automata. University of Illinois Press, Urbana, IL, USA, 1970. xxvi + 375 pp. LCCN QA267.5.S4 B87. A classic collection of papers on cellular automata intended as a companion to von Neumann’s Theory of Self Reproducing Automata [vNB66]. Includes papers by Burks, Thatcher, Moore, Myhill, Ulam, and Holland. See [Hol62].

Campbell:1987:NSP

[Cam87] David K. Campbell. Nonlinear science — from paradigms to practicalities. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):218–262, 1987. CODEN LASCDI. ISSN 0273-7116.

URL http://library.lanl.gov/cgi-bin/getfile?15-17.pdf.

Carlson:2003:HUS

[Car03] Bengt Carlson. How Ulam set the stage. Bulletin of the Atomic Scientists, 59(4):46–51, July 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chung:1979:MDT

[CEG^{+}79] F. R. K. Chung, P. Erdős, R. L. Graham, S. M. Ulam, and F. F. Yao. Minimal decompositions of two graphs into pairwise isomorphic subgraphs. In Proceedings of the Tenth Southeastern Conference on Combinatorics, Graph Theory and Computing (Florida Atlantic University, Boca Raton, Florida, April 2–6, 1979), Congress. Numer., XXIII–XXIV, pages 3–18. Utilitas Mathematica Publishing Inc., Winnipeg, Manitoba, Canada, 1979.

Cooper:1989:CCR

[CES89] Necia Grant Cooper, Roger Eckhardt, and Nancy Shera, editors. From cardinals to chaos: reflections on the life and legacy of Stanisław Ulam. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36494-9, 0-521-36734-4 (paperback). 318 + 2 pp. LCCN QA29.U4 F76 1989.

Capobianco:1971:RTG

[CFK71] Michael Capobianco, J. B. Frechen, and M. Krolik, editors. Recent trends in graph theory: proceedings of the first New York City graph theory conference held on June 11, 12, and 13, 1970, volume 215 of Lecture Notes in Mathematics. Springer-Verlag Inc., New York, NY, USA, 1971. ISBN 3-540-05386-7 (Berlin), 0-387-05386-7 (New York). LCCN ????

Cipra:2000:BCE

[Cip00] Barry A. Cipra. The best of the 20th Century: Editors name top 10 algorithms. SIAM News, 33(4):1–2, May 2000. ISSN 0036-1437.

URL https://archive.siam.org/pdf/news/637.pdf.

Ciesielski:2009:SUH

[CR09] Krzysztof Ciesielski and Themistocles M. Rassias. On Stan Ulam and his mathematics. The Australian Journal of Mathematical Analysis and Applications (AJMAA), 6(1):1–9, 2009. ISSN

1449-5910. URL http://ajmaa.org/searchroot/files/pdf/
v6n1/v6i1p1.pdf;   http://ajmaa.org/Volumes/Volume%206%

20Issue%201%202009.php. Special Issue in Honor of the 100th Anniversary of S. M. Ulam.

Cohen:1944:ASV

[CU44] L. Cohen and Stanisław M. Ulam. On the algebra of systems of vectors and some problems in kinematics. Bulletin of the American Mathematical Society, 50(??):61–??, ???? 1944. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Caillois:1966:RIG

[CU^{+}66] Roger Caillois, Stanisław M. Ulam, et al., editors. Rencontres Internationales de Genève “Le robot, la bete et l’homme” (1965). Éditions de la Baconnière, Neuchâtel, Switzerland, 1966. LCCN ????

Dauxois:2008:FPU

[Dau08] Thierry Dauxois. Fermi, Pasta, Ulam and a mysterious lady. arxiv.org, ??(??):3, January 10, 2008. URL http://arxiv.org/

abs/0801.1590.

Doolen:1985:PCT

[DKU85] Gary D. Doolen, Mark Kac, and Stanisław M. Ulam, editors. Proceedings of the Conference on Transport and Propagation in Nonlinear Systems, held at the Los Alamos National Laboratory, Los Alamos, New Mexico, May 1984, volume 39(5/6) of Journal of statistical physics, 0022-4715. Plenum Press, New York, NY, USA; London, UK, June 1985. LCCN QC175 .J826.

Dauxois:2005:FPU

[DPR05] Thierry Dauxois, Michel Peyrard, and Stefano Ruffo. The Fermi– Pasta–Ulam ‘numerical experiment’: history and pedagogical perspectives. European Journal of Physics, 26(5):S3–S11, 2005. CO-DEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

URL https://arxiv.org/pdf/nlin/0501053.pdf.

Dubovickii:1976:SPU

[Dub76] A. J. Dubovickii. Solution of a problem of S. Ulam on optimal matching of segments. Mathematics of the USSR — Izvestiya, 10 (3):639, 1976. CODEN MUSIAE. ISSN 0025-5726 (print), 2169-

5075 (electronic). URL http://stacks.iop.org/0025-5726/10/
i=3/a=A09.

Dyson:2002:POT

[Dys02] George Dyson. Project Orion: the true story of the atomic spaceship. Henry Holt and Company, New York, NY, USA, 2002. ISBN 0-8050-5985-7. xv + 345 pp. LCCN TL783.5 .D95 2002. URL

http://www.loc.gov/catdir/bios/hol051/2001046500.html;
 http://www.loc.gov/catdir/description/hol021/2001046500.
html.

Dyson:2012:MC

[Dys12a] George Dyson. Monte Carlo. In Turing’s cathedral: the origins of the digital universe [Dys12b], chapter 10, pages 175– 199. ISBN 0-375-42277-3 (hardcover), 1-4000-7599-8 (paperback). LCCN QA76.17 .D97 2012. Pages 191–192 describe the origin of the Monte Carlo method.

Dyson:2012:TCO

[Dys12b] George Dyson. Turing’s cathedral: the origins of the digital universe. Pantheon Books, New York, NY, USA, 2012. ISBN 0-375-42277-3 (hardcover), 1-4000-7599-8 (paperback). xxii + 401 pp. LCCN QA76.17 .D97 2012.

Dyson:2012:UD

[Dys12c] George Dyson. Ulam’s demons. In Turing’s cathedral: the origins of the digital universe [Dys12b], chapter 11, pages 200– 224. ISBN 0-375-42277-3 (hardcover), 1-4000-7599-8 (paperback). LCCN QA76.17 .D97 2012. Pages 191–192 describe the origin of the Monte Carlo method.

Eckhardt:1987:SUJ

[Eck87] Roger Eckhardt. Stan Ulam, John von Neumann, and the Monte Carlo method. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):131–137, 1987. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?
00326867.pdf;   http://library.lanl.gov/cgi-bin/getfile?

15-13.pdf. With contributions by Tony Warnock, Gary D. Doolen and John Hendricks, Stanisław Ulam 1909–1984.

Erdos:1944:SCP

[EU44a] P. Erdős and Stanisław M. Ulam. Some combinatorial problems in set theory. Preliminary report. Bulletin of the American Mathematical Society, 50(??):57–??, ???? 1944. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1944:OG

[EU44b] Cornelius J. Everett and S. Ulam. On ordered groups. Bulletin of the American Mathematical Society, 50(??):496–??, ???? 1944. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1945:OG

[EU45a] Cornelius J. Everett and S. Ulam. On ordered groups. Transactions of the American Mathematical Society, 57:208–216, 1945. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Everett:1945:PA

[EU45b] Cornelius J. Everett and S. Ulam. Projective algebra. I. Bulletin of the American Mathematical Society, 51(??):59–??, ???? 1945. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1946:PA

[EU46] Cornelius J. Everett and S. Ulam. Projective algebra. I. American Journal of Mathematics, 68:77–88, 1946. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Everett:1948:MS

[EU48a] Cornelius J. Everett and S. Ulam. Multiplicative systems. I. Proceedings of the National Academy of Sciences of the United States of America, 34(8):403–405, August 1948. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Everett:1948:PDM

[EU48b] Cornelius J. Everett and S. Ulam. On the problem of determination of mathematical structures by their endomorphisms. Bulletin of the American Mathematical Society, 54(??):646–??, ???? 1948. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1949:MSa

[EU49a] Cornelius J. Everett and S. M. Ulam. Multiplicative systems, I. Bulletin of the American Mathematical Society, 55(??):51–??, ???? 1949. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1949:MSb

[EU49b] Cornelius J. Everett and S. M. Ulam. Multiplicative systems, II. Bulletin of the American Mathematical Society, 55(??):51–??, ???? 1949. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1949:MSc

[EU49c] Cornelius J. Everett and S. M. Ulam. Multiplicative systems, III. Bulletin of the American Mathematical Society, 55(??):51–52, ???? 1949. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1950:CTR

[EU50a] C. J. Everett and S. Ulam. Considerations on thermonuclear reactions. Classified report LA-1158, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, June 16, 1950. 36 pp.

Everett:1950:ILMb

[EU50b] C. J. Everett and S. Ulam. Ignition of a large mass of deuterium by a burning D–T mixture: Problem II. Classified report LA-1124, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, June 16, 1950. 22 pp.

Everett:1950:ILMa

[EU50c] C. J. Everett and S. Ulam. Ignition of a large mass of deuterium by a burning deuterium–tritium mixture: Problem I. Classified report LA-1076, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, March 7, 1950. 40 pp.

Everett:1950:ACB

[EU50d] Cornelius J. Everett and Stanisław M. Ulam. On an application of a correspondence between matrices over real algebras and matrices of positive real numbers. Bulletin of the American Mathematical Society, 56(??):63–??, ???? 1950. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1950:RWH

[EU50e] Cornelius J. Everett and Stanisław M. Ulam. Random walk and the Hamilton–Jacobi equation. Bulletin of the American Mathematical Society, 56(??):63–64, ???? 1950. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Everett:1955:MPP

[EU55] Cornelius J. Everett and Stanisław M. Ulam. On a method of propulsion of projectiles by means of external nuclear explosions. Technical Report LAMS-1955, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1955.

Everett:1961:MCC

[EU61] Cornelius J. Everett and Stanisław M. Ulam. Nuclear propelled vehicle, such as a rocket. British Patent 877,392., 1961.

Everett:1966:SPG

[EU66] Cornelius J. Everett and S. M. Ulam. On some possibilities of generalizing the Lorentz group in the Special Relativity Theory. Journal of Combinatorial Theory, 1:248–270, 1966. CODEN JCTHAR. ISSN 0021-9800 (print), 1878-1756 (electronic).

Erdos:1968:ESU

[EU68] Paul Erdős and S. Ulam. On equations with sets as unknowns. Proceedings of the National Academy of Sciences of the United States of America, 60(4):1189–1195, 1968. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Everett:1969:EIP

[EU69] Cornelius J. Everett and Stanisław M. Ulam. The entropy of interacting populations. Technical Report LA-4256, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1969.

Erdos:1971:SPR

[EU71] P. Erdős and S. Ulam. Some probabilistic remarks on Fermat’s last theorem. Rocky Mountain Journal of Mathematics, 1(4):613–616, Fall 1971. CODEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic). URL http://projecteuclid.org/

euclid.rmjm/1250131729.

Fleming:1969:IME

[FB69] Donald Fleming and Bernard Bailyn, editors. The intellectual migration: Europe and America, 1930–1960. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 1969. ISBN 0-674-33411-6, 0-674-33412-4. 748 pp. LCCN E169.1 .F6.

Finkel:1985:USW

[FE85] L. H. Finkel and G. M. Edelman. Stanisław Ulam — the warmth and brilliance of an eclectic mind. Letters in Mathematical Physics, 10(2–3):243–245, ???? 1985. CODEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Feigenbaum:1985:ISU

[Fei85] Mitchell Feigenbaum. An interview with Stan Ulam and Mark Kac. Journal of Statistical Physics, 39(5–6):455–476, June 1985. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01008345. Reprinted from Los Alamos Science, Fall, 1982.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Flato:1984:SMU

[Fla84] Moshé Flato. Stanisław Marcin Ulam (April 13, 1909–May 13, 1984). Letters in Mathematical Physics, 8(4):257, ???? 1984. CO-DEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Flato:1985:MUS

[Fla85] Moshé Flato. In memory of Stanisław Ulam. Letters in Mathematical Physics, 10(2–3):95, ???? 1985. CODEN LMPHDY. ISSN 0377-9017 (print), 1573-0530 (electronic).

Ford:2015:BHB

[For15] Kenneth William Ford. Building the H bomb: a personal history. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2015. ISBN 981-4632-07-4 (hardcover), 981-4618-79-9 (paperback). 240 (est.) pp. LCCN UG1282.A8 F67

2015. URL http://www.worldscientific.com/worldscibooks/
10.1142/9269.

Fermi:1955:SNP

[FPU55] Enrico Fermi, John Pasta, and Stanisław M. Ulam. Studies of nonlinear problems. I. Technical Report LA-1940, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, May 1955. 22

pp. URL http://www.osti.gov/accomplishments/documents/

fullText/ACC0041.pdf. Also in Enrico Fermi: Collected Papers, volume 2, edited by Edoardo Amaldi, Herbert L. Anderson, Enrico Persico, Emilio Segré, and Albedo Wattenberg. Chicago: University of Chicago Press, 1965, pages 978–988.

Froyland:1999:UMR

[Fro99a] Gary Froyland. Ulam’s method for random interval maps. Nonlinearity (Bristol), 12(4):1029, 1999. CODEN NONLE5. ISSN 0951-7715 (print), 1361-6544 (electronic). URL http://stacks.

iop.org/0951-7715/12/i=4/a=318.

Froyland:1999:UUM

[Fro99b] Gary Froyland. Using Ulam’s method to calculate entropy and other dynamical invariants. Nonlinearity (Bristol), 12(1):79, 1999. CODEN NONLE5. ISSN 0951-7715 (print), 1361-6544 (electronic).

URL http://stacks.iop.org/0951-7715/12/i=1/a=006.

Frahm:2010:UMC

[FS10] K. M. Frahm and D. L. Shepelyansky. Ulam method for the Chirikov standard map. European Physical Journal B: Condensed Matter and Complex Systems, 76(1):57–68, July 2010. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic).

Gamow:1970:RSC

[Gam70a] G. Gamow. My world line: An informal autobiography. In My World Line: An Informal Autobiography. Viking Press, New York, NY, USA, 1970. LCCN ???? Foreword by Stanisław M. Ulam.

Gamow:1970:MWL

[Gam70b] George Gamow. My World Line: An Informal Autobiography. Viking, New York, NY, USA, 1970. 178 pp. LCCN ???? URL

http://adsabs.harvard.edu/abs/1970mwla.book.....G. Fore-

word by Stanisław M. Ulam.

Gardner:2001:CBM

[Gar01a] Martin Gardner. The colossal book of mathematics: classic puzzles, paradoxes, and problems: number theory, algebra, geometry, probability, topology, game theory, infinity, and other topics of recreational mathematics. W. W. Norton & Co., New York, NY, USA, 2001. ISBN 0-393-02023-1. xi + 724 pp. LCCN QA95 .G245 2001.

Gardner:2001:PS

[Gar01b] Martin Gardner. Packing spheres. In The colossal book of mathematics: classic puzzles, paradoxes, and problems: number theory, algebra, geometry, probability, topology, game theory, infinity, and other topics of recreational mathematics [Gar01a], chapter 10, pages 128–136. ISBN 0-393-02023-1. LCCN QA95 .G245 2001. Page 135 mentions a conjecture by Stan Ulam in 1972 that “the maximal density (≈ 0.74048) for packing congruent spheres is smaller than that for any other convex body” [LZ12, page 1547].

Gleason:2002:AUA

[Gle02] Abbott Gleason. Adam Ulam: 8 April 1922–28 March 2000. Proceedings of the American Philosophical Society, 146(4):415–418, December 2002. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
1558317.

Gardiner:1956:CSI

[GLMU56] Verna Gardiner, R. Lazarus, N. Metropolis, and S. Ulam. On certain sequences of integers defined by sieves. Mathematics Magazine, 29:117–122, 1956. CODEN MAMGA8. ISSN 0011-801x.

Gelernter:1961:BRM

[GM61] H. L. Gelernter and Benoît Mandelbrot. Book review: S. M. Ulam, A Collection of Mathematical Problems, (Tracts in Pure and Applied Mathematics, Number 8) $5.00, 1960 Interscience, New York 150. Information and Control, 4(1):81–82, March 1961. CODEN IFCNA4. ISSN 0019-9958 (print), 1878-2981

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0019995861800405.

Goad:1987:SAC

[Goa87] Walter B. Goad. Sequence analysis — contributions by Ulam to molecular genetics. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):288–291, 1987. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?
15-21.pdf.

Goldberger:1999:AEF

[Gol99] Marvin L. Goldberger. In appreciation: Enrico Fermi (1901– 1954): The complete physicist. Physics in Perspective (PIP), 1 (3):328–336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.com/

article/10.1007/s000160050024;   http://www.springerlink.
com/content/y1h02gr3rv7wurrl/.

Goncharov:1996:TMA

[Gon96a] German A. Goncharov. Thermonuclear milestones: (1) the American effort. Physics Today, 49(11):45–48, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMR

[Gon96b] German A. Goncharov. Thermonuclear milestones: (3) the race accelerates. Physics Today, 49(11):56–61, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goodchild:2004:ETRb

[Goo04] Peter Goodchild. Edward Teller, the real Dr. Strangelove. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01669-6. xxv + 469 + 16 pp. LCCN QC16.T37 G66 2004. URL http://

alsos.wlu.edu/information.aspx?id=2237.

Graham:1987:SMG

[Gra87] Ronald L. Graham. A similarity measure for graphs — reflections on a theme of Ulam. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):114–121, 1987. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?
15-10.pdf.

Hansen:1992:LUS

[Han92] Chuck Hansen. Letter: Ulam’s shock waves. Bulletin of the Atomic Scientists, 48(1):45, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BHM91].

Hansen:1997:LMF

[Han97] Chuck Hansen. Letter: More fallout on fallout. Bulletin of the Atomic Scientists, 53(5):3, 58, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM97].

Harlow:1987:EWN

[Har87] Francis H. Harlow. Early work in numerical hydrodynamics. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):144, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://library.

lanl.gov/cgi-bin/getfile?15-14.pdf.

Hargittai:2006:MSF

[Har06] István Hargittai. The Martians of Science: five physicists who changed the twentieth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517845-9 (hardcover). xxiv + 313 + 32 pp. LCCN QC15 .H27 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005029427-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2005029427-b.html;   http://www.loc.gov/catdir/toc/ecip061/
2005029427.html.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hasslacher:1987:DF

[Has87] Brosl Hasslacher. Discrete fluids. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):175–217, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-16.pdf.

Hawkins:1987:SPM

[Haw87] David Hawkins. The spirit of play — A memoir for Stan Ulam. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984): 39–51, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://

library.lanl.gov/cgi-bin/getfile?15-06.pdf.

Hill:2018:SSS

[Hil18] Theodore Hill. Slicing sandwiches, states, and solar systems. American Scientist, 106(1):42, 2018. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Hirsch:1990:HBW

[HM90] Daniel Hirsch and William G. Mathews. The H-bomb: Who really gave away the secret? Bulletin of the Atomic Scientists, 46(1):22– 30, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kno90, MHM90, BHM91, HM97].

Hirsch:1997:LHB

[HM97] Daniel Hirsch and William G. Mathews. Letter: H-bomb secrets. Bulletin of the Atomic Scientists, 53(4):3, 58, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90, KAS96] and comment [Han97].

Hoare:2002:SU

[Hoa02] G. T. Q. Hoare. Stanisław Ulam 1909–1984. Mathematics Today, 38(2):148–154, ???? 2002. ISSN 1361-2042.

Holland:1962:OLT

[Hol62] John H. Holland. Outline for a logical theory of adaptive systems. Journal of the ACM, 9(3):297–314, July 1962. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Early work on genetic algorithms. Reprinted in [Bur70].

Haigh:2014:ABE

[HPR14] T. Haigh, M. Priestley, and C. Rope. Los Alamos bets on ENIAC: Nuclear Monte Carlo simulations, 1947–1948. IEEE Annals of the History of Computing, 36(3):42–63, July 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hyers:1941:AIP

[HU41] D. H. Hyers and Stanisław M. Ulam. On approximate isometries. Preliminary report. Bulletin of the American Mathematical Society, 47(??):708–??, ???? 1941. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Hyers:1942:AIS

[HU42] D. H. Hyers and Stanisław M. Ulam. Approximate isometries of the space of continuous functions. Bulletin of the American Mathematical Society, 48(??):368–??, ???? 1942. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Hyers:1945:AI

[HU45] D. H. Hyers and S. M. Ulam. On approximate isometries. Bulletin of the American Mathematical Society, 51(4):288–292, 1945. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

URL http://projecteuclid.org/euclid.bams/1183506880.

Hyers:1947:AIS

[HU47] D. H. Hyers and S. M. Ulam. Approximate isometries of the space of continuous functions. Annals of Mathematics (2), 48:285–289, 1947. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Hyers:1951:ACF

[HU51] D. H. Hyers and S. M. Ulam. Approximately convex functions. Bulletin of the American Mathematical Society, 59(??):300–301, ???? 1951. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Hyers:1952:ACF

[HU52] D. H. Hyers and S. M. Ulam. Approximately convex functions. Proceedings of the American Mathematical Society, 3:821–828, 1952. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Holladay:1960:SCP

[HU60] J. C. Holladay and Stanisław M. Ulam. On some combinatorial problems in patterns of growth, I. Notices of the American Mathematical Society, 7(??):234–??, ???? 1960. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Huang:1976:POW

[Hua76] Kerson Huang, editor. Physics and our world: a symposium in honor of Victor F. Weisskopf, Massachusetts Institute of Technology, 1974, volume 28 of AIP conference proceedings. American Institute of Physics, New York, NY, USA, 1976. ISBN 0-88318-127-4. LCCN QC28 .P48.

Hurd:1985:NEM

[Hur85] Cuthbert C. Hurd. A note on early Monte Carlo computations and scientific meetings. Annals of the History of Computing, 7(2):141– 155, April/June 1985. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1985/pdf/a2141.pdf;
 http://www.computer.org/annals/an1985/a2141abs.htm. In-

cludes typeset reprint of [RUvN47].

Ji:2013:CAM

[JML13] Hao Ji, Michael Mascagni, and Yaohang Li. Convergence analysis of Markov Chain Monte Carlo linear solvers using Ulam–von Neumann algorithm. SIAM Journal on Numerical Analysis, 51(4): 2107–2122, ???? 2013. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Jung:2001:HUR

[Jun01] Soon-Mo Jung. Hyers–Ulam–Rassias stability of functional equations in mathematical analysis. Hadronic Press, Inc., Palm Harbor, FL, USA, 2001. ISBN 1-57485-051-2. ix + 256 + 2 pp.

Karpinski:1977:FCT

[Kar77] Marek Karpinski, editor. Fundamentals of computation theory: proceedings of the 1977 International FCT-Conference, Poznan-Kornik, Poland, September 19–23, 1977, volume 56 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1977. CODEN LNCSD9. ISBN 0-387-08442-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1,L43

no.56. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t0056.htm;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=56.

Khariton:1996:WIW

[KAS96] Yuli Khariton, Viktor Adamskii, and Yuri Smirnov. The way it was. Bulletin of the Atomic Scientists, 52(9):54–59, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.highbeam.com/doc/

1G1-18855733.html. See comment [HM97].

Kemeny:1970:RML

[Kem70] J. G. Kemeny. Reviews: Mathematics and Logic, Retrospect and Prospects, by Mark Kac and Stanislaw M. Ulam. American Mathematical Monthly, 77(3):316, March 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knoll:1990:LHB

[Kno90] Erwin Knoll. Letter: H-bomb secret never was. Bulletin of the Atomic Scientists, 46(3):53–54, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90].

Kister:1957:EC

[KSU^{+}57] J. Kister, P. Stein, S. Ulam, W. Walden, and M. Wells. Experiments in chess. Journal of the ACM, 4(2):174–177, April 1957. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Kac:1968:MLR

[KU68] Mark Kac and Stanisław M. Ulam. Mathematics and logic: Retrospect and prospects. Britannica perspective. Frederick A. Praeger, Inc., New York, NY, USA, 1968. ISBN 0-14-021253-1. x + 170 pp. LCCN QA93 .K113m; QA93 .K33; QA93 .K25. The text of this book first appeared as article with the same title in Britannica Perspectives, volume 1, 557732. Chicago: Encyclopaedia Britannica, Inc., 1968. Translated into French (1973), into Serbo-Croatian (1977), and into Spanish (1979).

Kac:1971:MLR

[KU71] M. Kac and S. Ulam. Matematika i logika. Retrospektiva i perspektivy. Izdat. “Mir”, Moscow, 1971. 251 pp. Translated from the English by N. I. Plužnikova, Edited by I. M. Jaglom, Contemporary Mathematics. Popular series.

Kac:1973:MLR

[KU73] Mark Kac and Stanisław M. Ulam. Mathématiques et logique: rétrospective et perspectives. Dunod, Paris, France, 1973. 178 pp. LCCN QA93 .K2514 1973.

Kac:1979:ML

[KU79] Mark Kac and Stanisław M. Ulam. Mathematics and logic. Penguin Books, Harmondsworth, UK, 1979. ISBN 0-14-080387-4. 204 pp. Retrospect and prospects, Abridged reprint of the 1968 edition.

Kac:1992:ML

[KU92] Mark Kac and Stanisław M. Ulam. Mathematics and logic. Dover Publications, Inc., New York, NY, USA, 1992. ISBN 0-486-67085-6. x + 170 pp. Reprint of the 1968 original.

Kuratowski:1980:HCP

[Kur80] Kazimierz Kuratowski. A half century of Polish mathematics: remembrances and reflections, volume 108 of International series in pure and applied mathematics. Państwowe Wydawnictwo Naukowe (distributed by Ars Polona), Warszawa, Poland, 1980. ISBN 83-01-00819-9. viii + 204 pp. LCCN QA27.P67 K7813. Preface by Stanisław M. Ulam.

Kraft:1964:MCM

[KW64] Rose Kraft and Carl J. Wensrich. Monte Carlo methods: a bibliography covering the period 1949 to 1963. Report UCRL-7823 and UC-32 and TID-4500 (32nd edition), University of California Radiation Laboratory, Berkeley, CA, USA, April 1, 1964. iii + 110

pp. URL https://catalog.hathitrust.org/Record/007292355.

Supersedes report UCRL-658.

Lomnicki:1933:GGZ

[LU33a] Z. Lomnicki and Stanisław M. Ulam. Über die Gesetze der grossen Zahlen. (German) [On the laws of large numbers]. Annales de la Société Polonaise de Mathématiques, 12(??):118–??, ???? 1933.

Lomnicki:1933:GW

[LU33b] Z. Lomnicki and Stanisław M. Ulam. Über die Grundlagen der Wahrscheinlichkeitsrechnung. (German) [On the basic of probability calculation]. Annales de la Société Polonaise de Mathématiques, 12(??):115–??, ???? 1933.

Lomnicki:1934:TME

[LU34] Z. Lomnicki and S. Ulam. Sur la théorie de la mesure dans les espaces combinatoires et son application au calcul des probabiliteés. I. variables independantes. (French) [On measure theory in combinatorial spaces and its application to the computation of probabilities]. Fundamenta mathematicae, 23:237–278, 1934. ISSN 0016-2736 (print), 1730-6329 (electronic).

Lagarias:2012:MPR

[LZ12] Jeffrey C. Lagarias and Chuanming Zong. Mysteries in packing regular tetrahedra. Notices of the American Mathematical Society, 59(11):1540–1549, December 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.ams.

org/notices/201211/rtx121101540p.pdf.

Macrae:1992:JNSa

[Mac92] Norman Macrae. John von Neumann: The Scientific Genius Who Pioneered the Modern Computer, Game Theory, Nuclear Deterrence, and Much More. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-41308-1. x + 405 pp. LCCN QA29.V66 M33 1992. US$25.00 (US$31.50 Can.).

Mark:1987:SUA

[Mar87] Carson Mark. Stan Ulam: From above the fray. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):33, 1987. CO-DEN LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/

cgi-bin/getfile?15-04.pdf.

Mauldin:1981:SBM

[Mau81] R. Daniel Mauldin, editor. The Scottish book: mathematics from the Scottish Cafe. Birkhäuser Boston Inc., Cambridge, MA, USA, 1981. ISBN 3-7643-3045-7. LCCN QA43 .S39. First edition 1957 and 1977 by Stanisław Ulam and distributed by the Los Alamos Laboratory. Contains mathematical problems, collected in a notebook by Polish mathematicians in the Scottish coffee house in Lwów from 1935 to 1941. Includes selected papers from the Scottish Book Conference, held at North Texas State University in May 1979.

Mauldin:1987:PNS

[Mau87] R. Daniel Mauldin. Probability and nonlinear systems. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):52–90, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://library.lanl.

gov/cgi-bin/getfile?15-07.pdf.

Mauldin:2015:SBM

[Mau15] R. Daniel Mauldin. The Scottish Book. Mathematics from The Scottish Café, with Selected Problems from The New Scottish Book. Springer International Publishing, Cham, Switzerland, second edition, December 2015. ISBN 3-319-22896-X (print), 3-319-22897-8 (e-book), 3-319-22898-6, 3-319-79432-9. xvii + 322 pp. LCCN

QA39.3; QA43 .S39 2015. URL https://link.springer.com/
book/10.1007/978-3-319-22897-6.

May:1969:BRB

[May69] Kenneth O. May. Book review: Mathematics and Logic. Retrospect and Prospect by Mark Kac; Stanislaw Ulam. Isis, 60(1):112–113, Spring 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/229029.

McMillan:2005:RJR

[McM05] Priscilla Johnson McMillan. The ruin of J. Robert Oppenheimer, and the birth of the modern arms race. Viking, New York, NY, USA, 2005. ISBN 0-670-03422-3. viii + 373 + 16 pp. LCCN QC16.O62 M36 2005.

Metropolis:1976:BRA

[Met76] N. Metropolis. Book review: Adventures of a Mathematician, by S. M. Ulam. Science (New Series), 193(4254):568– 569, August 1976. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://links.jstor.org/sici?sici=0036-8075%
2819760813%293%3A193%3A4253%3C568%3AM%3E2.0.CO%3B2-M.

Metropolis:1987:BMC

[Met87] N. Metropolis. The beginning of the Monte Carlo method. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):125– 130, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://

library.lanl.gov/cgi-bin/getfile?15-12.pdf.

Moss:1990:LWF

[MHM90] Norman Moss, Daniel Hirsch, and William G. Mathews. Letters: What Fuchs didn’t know, and when he didn’t know it. Bulletin of the Atomic Scientists, 46(4):51–52, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90].

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Moore:1982:BRBa

[Moo82] Gregory H. Moore. Book review: A half century of Polish mathematics: Remembrances and reflections: By Kazimierz Kuratowski. Preface by Stanislaw Ulam. Oxford (Pergamon) and Warsaw (PWN, Polish Scientific Publishers). 1980. viii + 200 pp. 72 photographs. Historia Mathematica, 9(4):489–495, November 1982. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086082901136.

Metropolis:1976:EDM

[MOR76a] N. (Nicholas) Metropolis, Steven A. Orszag, and Gian-Carlo Rota, editors. Essays dedicated to S. M. Ulam: proceedings of the first Los Alamos Symposium on Mathematics in the Natural Sciences, Surveys in applied mathematics. Academic Press, New York, NY, USA, 1976. ISBN 0-12-492150-7. LCCN QA36 .L896 1974; QA1 .L588 1974.

Metropolis:1976:SAM

[MOR76b] N. (Nicholas) Metropolis, Steven A. Orszag, and Gian-Carlo Rota, editors. Surveys in applied mathematics: essays dedicated to S. M. Ulam: proceedings. Academic Press, New York, NY, USA, 1976. LCCN QA1.L588 1974.

Menzel:1959:QTP

[MSU59] M. T. Menzel, P. R. Stein, and Stanisław M. Ulam. Quadratic transformations. Part I. Technical Report LA-2305, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1959.

Mazur:1930:UAG

[MU30] S. Mazur and Stanisław M. Ulam. Über unendliche Abelsche Gruppen. (German) [On infinite Abelian groups]. Annales de la Société Polonaise de Mathématiques, 9(??):204–??, ???? 1930.

Mazur:1931:IAN

[MU31] S. Mazur and Stanisław M. Ulam. Über isometrische Abbildungen von normierten Vektorraumen. (German) [On isometric transformations of normed vector spaces]. Annales de la Société Polonaise de Mathématiques, 10(??):127–??, ???? 1931.

Mazur:1932:TIE

[MU32] S. Mazur and Stanisław M. Ulam. Sur les transformations isometriques d’espaces vectoriels normés. Comptes Rendus de l’Academie des Sciences, 194(??):946–948, ???? 1932.

Metropolis:1949:MSM

[MU49a] N. C. Metropolis and Stanisław M. Ulam. On motions of systems of mass points randomly distributed on the infinite line. Bulletin of the American Mathematical Society, 55(??):670–671, ???? 1949. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Metropolis:1949:MCM

[MU49b] Nicholas Metropolis and S. Ulam. The Monte Carlo method. Journal of the American Statistical Association, 44(247):335–341, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://links.jstor.org/sici?
sici=0162-1459%28194909%2944%3A247%3C335%3ATMCM%3E2.0.
CO%3B2-3;   http://www.jstor.org/stable/2280232.

Metropolis:1953:PRA

[MU53] N. Metropolis and S. Ulam. A property of randomness of an arithmetical function. American Mathematical Monthly, 60(4):252–253, April 1953. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://links.jstor.org/sici?sici=0002-
9890%28195304%2960%3A4%3C252%3AAPOROA%3E2.0.CO%3B2-T.

Meisters:1967:VHS

[MU67] G. H. Meisters and S. M. Ulam. On visual hulls of sets. Proceedings of the National Academy of Sciences of the United States of America, 57(5):1172–1174, 1967. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). May.

Mycielski:1969:PPN

[MU69] Jan Mycielski and S. M. Ulam. On the pairing process and the notion of genealogical distance. Journal of Combinatorial Theory, 6:227–234, 1969. CODEN JCTHAR. ISSN 0021-9800 (print), 1878-1756 (electronic).

Mark:1987:MM

[MU87a] J. Carson Mark and S. M. Ulam. A memorable memo. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):294, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://library.lanl.

gov/cgi-bin/getfile?15-23.pdf. Stanisław Ulam 1909–1984.

Mauldin:1987:MPG

[MU87b] R. Daniel Mauldin and S. M. Ulam. Mathematical problems and games. Advances in Applied Mathematics, 8(3):281–344, 1987. ISSN 0196-8858 (print), 1090-2074 (electronic).

Metropolis:19xx:PRA

[MUxx] N. Metropolis and S. Ulam. A property of randomness of an arithmetical function. Report AECU-2038; LADC-1177, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 19xx.

3 pp. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=4404316&query_id=0.

Mycielski:1977:CUI

[Myc77] Jan Mycielski. A conjecture of Ulam on the invariance of measure in Hilbert’s cube. Studia Mathematica, 60(1):1–10, 1977. CODEN SMATAZ. ISSN 0039-3223 (print), 1730-6337 (electronic).

Mycielski:1987:LUM

[Myc87] Jan Mycielski. Learning from Ulam: Measurable cardinals, ergodicity, and biomathematics. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):107–113, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-09.pdf.

Mycielski:1990:SMU

[Myc90] Jan Mycielski. Stanisław Marcin Ulam (1909–1984). Wiadom. Mat., 29(1):21–37, 1990. CODEN RPTWAD. ISSN 0373-8302.

Mysels:1990:WUS

[Mys90] K. J. Mysels. Was Stan Ulam lost in the Lost Cafe? Physics Today, 43(7):13, July 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Neyman:1974:HCT

[Ney74] Jerzy Neyman. The heritage of Copernicus: theories “pleasing to the mind”. MIT Press, Cambridge, MA, USA, 1974. ISBN 0-262-14021-7. x + 542 pp. LCCN Q126.8 .N48.

Oxtoby:1938:ESF

[OU38a] J. C. Oxtoby and S. M. Ulam. On the equivalence of any set of first category to a set of measure zero. Fundam. Math., Warszawa, 31:201–206, 1938.

Oxtoby:1938:EMT

[OU38b] J. C. Oxtoby and Stanisław M. Ulam. Existence of metrically transitive transformations. Preliminary report. Bulletin of the American Mathematical Society, 44(??):347–??, ???? 1938. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Oxtoby:1939:EMI

[OU39] J. C. Oxtoby and S. M. Ulam. On the existence of a measure invariant under a transformation. Annals of Mathematics (2), 40:

560–566, 1939. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Oxtoby:1941:MPH

[OU41] J. C. Oxtoby and S. M. Ulam. Measure-preserving homeomorphisms and metrical transitivity. Annals of Mathematics (2), 42: 874–920, 1941. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic).

Onorato:2015:RTF

[OVPL15] Miguel Onorato, Lara Vozella, Davide Proment, and Yuri V. Lvov. Route to thermalization in the α-fermi–pasta–ulam system. Proceedings of the National Academy of Sciences of the United States of America, 2015. CODEN PNASA6. ISSN 0027-8424 (print), 1091-

6490 (electronic). URL http://www.pnas.org/content/early/
2015/03/23/1404397112.abstract.

Patrascioiu:1987:EHC

[Pat87] Adrian Patrascioiu. The ergodic hypothesis: A complicated problem of mathematics and physics. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):263–279, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-18.pdf.

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN QC16.O62 P35 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005002173-d.
html;   http://www.loc.gov/catdir/enhancements/fy0723/
2005002173-b.html;   http://www.loc.gov/catdir/toc/ecip056/
2005002173.html.

Pasta:1959:HNW

[PU59] John R. Pasta and S. Ulam. Heuristic numerical work in some problems of hydrodynamics. Mathematical Tables and Other Aids to Computation, 13(65):1–12, January 1959. CODEN MTTCAS. ISSN 0891-6837.

Pasta:1965:SNP

[PUF65] J. R. Pasta, S. M. Ulam, and E. Fermi. Studies on nonlinear problems. In Collected works of Enrico Fermi, volume 2, page 978. Univ. of Chicago Press, Chicago, 1965.

Porter:2009:FPU

[PZHC09] Mason A. Porter, Norman J. Zabusky, Bambi Hu, and David K. Campbell. Fermi, Pasta, Ulam and the birth of experimental mathematics: a numerical experiment that Enrico Fermi, John Pasta, and Stanisław Ulam reported 54 years ago continues to inspire discovery. American Scientist, 97(3):214–221, May/June 2009. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.americanscientist.org/issues/
feature/2009/3/fermi-pasta-ulam-and-the-birth-of-experimental-

mathematics. The work referred to in the title is [FPU55].

Rassias:2010:FED

[Ras10] John Michael Rassias, editor. Functional equations, difference inequalities, and Ulam stability notions (F.U.N.). Mathematics research developments series. Nova Science Publishers, New York, NY, USA, 2010. ISBN 1-60876-461-3 (hardcover). ix + 211 pp. LCCN QA431 .F7886 2010.

Rassias:2012:FEM

[RB12] Themistocles M. Rassias and Janusz Brzdȩk, editors. Functional Equations in Mathematical Analysis, volume 52 of Springer optimization and its applications. Springer-Verlag Inc., New York, NY, USA, 2012. ISBN 1-4614-0054-6, 1-4614-0055-4 (e-book). ISSN 1931-6828. xvii + 748 pp. LCCN QA431 .F86 2012.

Redei:2005:JNS

[Réd05] Miklós Rédei. John von Neumann selected letters, volume 27 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1 (hardcover). ISSN 0899-2428. xxv + 301 pp. LCCN QA29.V66 A4 2005.

Reines:1972:CFOa

[Rei72] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Rider:1989:CCR

[Rid89] R. E. Rider. From cardinals to chaos — reflections on the life and legacy of Stanisław Ulam. Science, 246(4926):134, October 6, 1989. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Rota:1984:WSM

[Rot84] Gian-Carlo Rota. Words spoken at the memorial service for S. M. Ulam. The Mathematical Intelligencer, 6(4):40–42, 1984. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Rota:1985:BM

[Rot85] Gian-Carlo Rota. The barrier of meaning. Letters in Mathematical Physics, 10(2-3):97–99, 1985. CODEN LMPDHQ. ISSN 0377-9017 (print), 1573-0530 (electronic).

Rota:1986:MUS

[Rot86a] G. C. Rota. In memoriam of Stan Ulam — the barrier of meaning. Physica D, Nonlinear phenomena, 22(1–3):1–3, ???? 1986. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Rota:1986:PSM

[Rot86b] Gian-Carlo Rota, editor. Probability, statistical mechanics, and number theory: a volume dedicated to Mark Kac, volume 9 of Advances in mathematics. Supplementary studies. Academic Press, New York, NY, USA, 1986. ISBN 0-12-598543-6. xi + 194 pp. LCCN QA273 .P7954 1986.

Rota:1987:SUL

[Rot87] Gian-Carlo Rota. Stan Ulam: The lost cafe. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):23–32, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-03.pdf.

Rota:1989:BMM

[Rot89] Gian-Carlo Rota. The barrier of meaning. In memorium: Stanisław Ulam. Notices of the American Mathematical Society, 36(2):141– 143, 1989. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rota:1990:WUS

[Rot90] G. C. Rota. Was Stan Ulam lost in the Lost Cafe? — reply. Physics Today, 43(7):93, July 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Rassias:1982:SSP

[RR82] Themistocles M. Rassias and George M. Rassias, editors. Selected studies, physics-astrophysics, mathematics, history of science: a volume dedicated to the memory of Albert Einstein. North-Holland Publishing Co., Amsterdam, The Netherlands, 1982. ISBN 0-444-86161-0. xii + 392 pp. LCCN QB461 .S374 1982.

Richtmyer:1947:SMN

[RUvN47] Robert D. Richtmyer, Stanisław Ulam, and John von Neumann. Statistical methods in neutron diffusion. Technical Report LAMS-551, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, April 9, 1947. 22 pp. Republished in typeset form in [Hur85].

Redei:2012:HIP

[RW12] Miklós Rédei and Charlotte Werndl. On the history of the isomorphism problem of dynamical systems with special regard to von Neumann’s contribution. Archive for History of Exact Sciences, 66(1):71–93, January 2012. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=66&
issue=1&spage=71.

Stein:1987:IMS

[Ste87] Paul R. Stein. Iteration of maps, strange attractors, and number theory — an Ulamian potpourri. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):91–106, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-08.pdf.

Strogatz:1998:BRT

[Str98] Steven H. Strogatz. Book review: Thomas P. Weissert, The genesis of simulation in dynamics: Pursuing the Fermi–Pasta–Ulam problem. Computers in Physics, 12(6):574–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168754.

Schreier:1933:TCS

[SU33a] J. Schreier and S. Ulam. Sur les transformations continues des sphères euclidiennes. (French) [On continuous transformations of Euclidean spheres]. Comptes rendus de l’Académie des sciences, Paris, 197:967–968, 1933.

Schreier:1933:APNa

[SU33b] J. Schreier and S. Ulam. Über die Automorphismen der Permutationsgruppe der natürlichen Zahlenfolge. (German) [On the automorphism of permutation groups of natural number sequences].

Studia Mathematica, 4(??):134–141, ???? 1933. CODEN SMATAZ. ISSN 0039-3223 (print), 1730-6337 (electronic).

Schreier:1934:BTA

[SU34a] J. Schreier and S. Ulam. Über topologische Abbildungen der euklidischen Sphären. (German) [On topological transformations of the Euclidean sphere]. Fundamenta mathematicae, 23:102–118, 1934. ISSN 0016-2736 (print), 1730-6329 (electronic).

Schreier:1934:BIS

[SU34b] J. Schreier and Stanisław M. Ulam. Bemerkungen über die stetigen Abbildungen von Topologischen Raumen. (German) [Remark on constant transformations of topological spaces ]. Annales de la Société Polonaise de Mathématiques, 13(??):142–??, ???? 1934.

Schreier:1935:BGT

[SU35a] J. Schreier and S. Ulam. Eine Bemerkung über die Gruppe der topologischen Abbildungen der Kreislinie auf sich selbst. (German) [A remark n the group of topological transformations of the circle line on itself ]. Studia Mathematica, 5:155–159, 1935. CODEN SMATAZ. ISSN 0039-3223 (print), 1730-6337 (electronic).

Schreier:1935:NGG

[SU35b] J. Schreier and S. Ulam. Sur le nombre des générateurs d’un groupe topologique compact et connexe. (French) [On the number of generators of a topological compact and connected group]. Fundamenta mathematicae, 24:302–304, 1935. ISSN 0016-2736 (print), 1730-6329 (electronic).

Schreier:1937:APN

[SU37] J. Schreier and S. Ulam. Über die Automorphismen der Permutationsgruppe der natürlichen Zahlenfolge. (German) [On the automorphism of permutation groups of natural number sequences]. Fundamenta mathematicae, 28:258–260, 1937. ISSN 0016-2736 (print), 1730-6329 (electronic).

Stein:1955:SCC

[SU55] P. R. Stein and Stanisław M. Ulam. Study of certain combinatorial problems through experiments on computing machines. In ????, editor, Proceedings of the 1955 High-Speed Computer Conference (Louisiana State University, Baton Rouge, Louisiana, February 14– 16, 1955), pages 101–106. ????, ????, 1955. LCCN ????

Stein:1957:ECEa

[SU57a] P. R. Stein and Stanisław M. Ulam. Experiments in chess on electronic computing machines. Chess Review, ??(??):13–15, January 1957. Also in [SU57b].

Stein:1957:ECEb

[SU57b] P. R. Stein and Stanisław M. Ulam. Experiments in chess on electronic computing machines. Computers and Automation, ??(??):??, September 1957. CODEN CPAUAJ. ISSN 0010-4795, 0887-4549.

Stein:1959:CBR

[SU59] P. R. Stein and Stanisław M. Ulam. On certain binary reaction systems. Notices of the American Mathematical Society, 6(1):68– 69, February 1959. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Schrandt:1960:PGF

[SU60] R. G. Schrandt and S. M. Ulam. On patterns of growth of figures in two dimensions. Notices of the American Mathematical Society, 7(5):642, October 1960. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). Early studies of pattern development in cellular automata, including competitions between two kinds of patterns. Reprinted in [Bur70].

Stein:1963:NLT

[SU63] Paul R. Stein and Stanisław M. Ulam. Non-linear transformation studies on electronic computers. Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1963. 125 pp. LCCN QA601 .S82.

Stein:1964:NLT

[SU64] P. R. Stein and Stanisław M. Ulam. Non-linear transformation studies on electronic computers. Rozprawy Mat., 39:1–66, 1964. The Introduction and Part I are reprinted in [Bur70].

Schrandt:1967:RDG

[SU67a] R. G. Schrandt and Stanisław M. Ulam. On recursively defined geometrical objects and patterns of growth. Technical Report LA-3762, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1967. Reprinted in [Bur70].

Stein:1967:MNO

[SU67b] M. Stein and S. M. Ulam. Mathematical notes: An observation on the distribution of primes. American Mathematical Monthly, 74(1):43–44, 1967. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stein:1973:LNA

[SU73] P. R. Stein and S. Ulam. Lectures in nonlinear algebraic transformations. In A. O. Barut, editor, Studies in Mathematical Physics (Lectures Presented at the NATO Advanced Study Institute on Mathematical Physics held in Istanbul, August, 1970), pages 263–314. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973.

Sullivan:1984:SUT

[Sul84] Walter Sullivan. Stanisław Ulam, theorist on hydrogen bomb. New York Times, ??(??):??, May 15, 1984. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:

//www.nytimes.com/1984/05/15/obituaries/stanislaw-ulam-
theorist-on-hydrogen-bomb.html.

Stein:1964:MNV

[SUW64] M. L. Stein, S. M. Ulam, and M. B. Wells. Mathematical notes: A visual display of some properties of the distribution of primes. American Mathematical Monthly, 71(5):516–520, 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shepelyansky:2010:GMD

[SZ10] D. L. Shepelyansky and O. V. Zhirov. Google matrix, dynamical attractors, and Ulam networks. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 81 (3):036213:1–036213:9, March 2010. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://journals.aps.

org/pre/abstract/10.1103/PhysRevE.81.036213;   http://
link.aps.org/doi/10.1103/PhysRevE.81.036213.

Teller:1951:HDH

[TU51] Edward Teller and Stanisław Ulam. On heterocatalytic detonations. I. Hydrodynamic lenses and radiation mirrors. Report LAMS-1225, Los Alamos Scientific Laboratory of the University of California, Los Alamos, NM, USA, March 9, 1951.

20 pp. URL http://blog.nuclearsecrecy.com/wp-content/
uploads/2012/02/1951-Teller-Ulam-LAMS1225.pdf;   http:/
/www.nuclearnonproliferation.org/LAMS1225.pdf. This now-

classified report is believed to contain the Teller–Ulam hydrogenbomb trigger design. See [Har10, page 221], [Mac92, page 344], [Goo04, pages 179–181]. For a discussion of this report, see [For15, Chapter 1]. Copies found on the Web are heavily censored, with many pages completely blanked out.

Tarwater:1981:AMH

[TWHM81] Dalton Tarwater, John T. White, Carl Hall, and Marion E. Moore, editors. American mathematical heritage: algebra and applied mathematics[: papers of the third conference, held at the University of Texas at El Paso, November 1975, and the fourth conference, held at the University of Texas at Arlington, October 1976], volume 13 of Mathematics Series. Texas Technological University Press, Lubbock, TX, USA, 1981.

Ulam:1931:NTO

[UB31] Stanisław M. Ulam and Karol Borsuk. Über eine neue topologische Operation. (German) [On a new topological operation]. Annales de la Société Polonaise de Mathématiques, 10(??):125–126, ???? 1931.

Ulam:1964:FNE

[UdvB^{+}64] Stanisław M. Ulam, F. deLuzio, W. von Braun, M. Hunter, and I. Asimov. The future of nuclear energy in space: A panel discussion sponsored by the Aerospace Division, American Nuclear Society, at the 1963 winter meeting in New York City, NY on November 20, 1963. In R. F. Trapp, editor, ????, page ?? ????, ????, 1964. LCCN ???? Also in Nuclear News, July 1964.

Ulam:1948:MSSa

[UE48a] Stanisław M. Ulam and Cornelius J. Everett. Multiplicative systems in several variables. Part I. Technical Report LA-683, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1948.

Ulam:1948:MSSb

[UE48b] Stanisław M. Ulam and Cornelius J. Everett. Multiplicative systems in several variables. Part II. Technical Report LA-690, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1948.

Ulam:1948:MSSc

[UE48c] Stanisław M. Ulam and Cornelius J. Everett. Multiplicative systems in several variables. Part III. Technical Report LA-707, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1948.

Ulam:1944:TMP

[UH44] Stanisław M. Ulam and D. Hawkins. Theory of multiplicative processes. I. Technical Report LA-171, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1944.

Ulam:1954:SDE

[UH54] S. M. Ulam and D. H. Hyers. On the stability of differential expressions. Mathematics Magazine, 28:59–64, 1954. CODEN MAMGA8. ISSN 0011-801x.

Ulam:1932:QPT

[UK32] Stanisław M. Ulam and C. Kuratowski. Quelques propriétés topologiques du produit combinatoire. (French) [Some topological properties of the combinatorial product]. Fundamenta mathematicae, 19(??):247–251, ???? 1932. ISSN 0016-2736 (print), 1730-6329 (electronic).

Ulam:1933:CLT

[UK33] Stanisław M. Ulam and C. Kuratowski. Sur un coefficient lié aux transformations continues d’ensembles. (French) [On a coefficient related to continuous transformations of sets ]. Fundamenta mathematicae, 20(??):244–253, ???? 1933. ISSN 0016-2736 (print), 1730-6329 (electronic).

Ulam:1968:JN

[UKTS68] S. Ulam, H. W. Kuhn, A. W. Tucker, and Claude E. Shannon. John von Neumann, 1903–1957. Perspectives in American History, 2(??):235–269, ???? 1968. CODEN ???? ISSN 0079-0990.

Ulam:1969:JN

[UKTS69] Stanisław M. Ulam, H. W. Kuhn, A. W. Tucker, and Claude E. Shannon. John von Neumann, 1903–1957. In Fleming and Bailyn [FB69], pages 235–268. ISBN 0-674-33411-6, 0-674-33412-4. LCCN E169.1 .F6.

Ulam:1929:CFS

[Ula29a] Stanisław M. Ulam. Concerning functions of sets. Fundamenta mathematicae, 14(??):231–233, ???? 1929. ISSN 0016-2736 (print), 1730-6329 (electronic).

Ulam:1929:RGB

[Ula29b] Stanisław M. Ulam. Remark on the generalized Bernstein’s theorem. Fundamenta mathematicae, 13(??):281–283, ???? 1929. ISSN 0016-2736 (print), 1730-6329 (electronic).

Ulam:1930:BM

[Ula30a] Stanisław M. Ulam. Ein Betrag zum Massproblem. (German) [A contribution to the measure problem]. Annales de la Société Polonaise de Mathématiques, 9(??):198–??, ???? 1930.

Ulam:1930:EMG

[Ula30b] Stanisław M. Ulam. Über die Eindeutigkeit des Masses von Gerardenmengen. (German) [Uniqueness of the Gerard measure]. Annales de la Société Polonaise de Mathématiques, 9(??):200–??, ???? 1930.

Ulam:1930:VAM

[Ula30c] Stanisław M. Ulam. Über vollstandig additive Massfunktionen in abstrakten Raumen. (German) [On complete additive measure functions in abstract spaces]. Annales de la Société Polonaise de Mathématiques, 9(??):195–??, ???? 1930.

Ulam:1930:MAM

[Ula30d] Stanisław M. Ulam. Zur Masstheorie in der allgemeinen Mengenlehre. (German) [On measure theory in general set theory]. Fundamenta mathematicae, 16(??):140–150, ???? 1930. ISSN 0016-2736 (print), 1730-6329 (electronic). Also in Mengenlehre, edited by U. Felgner, 223-233. Darmstadt: Wissenschaftliche Gesellschaft, 1979.

Ulam:1930:TF

[Ula30e] Stanisław M. Ulam. Zur Theorie des Fixpunktes. (German) [On the theory of fixed points]. Annales de la Société Polonaise de Mathématiques, 9(??):201–202, ???? 1930.

Ulam:1931:ESI

[Ula31] Stanisław M. Ulam. Einige Satze über Mengen II-er Kategorie. (German) [A proposition on quantities of the second category]. Annales de la Société Polonaise de Mathématiques, 10(??):123–124, ???? 1931.

Ulam:1932:MP

[Ula32] Stanisław M. Ulam. Zum Massbegriff in Produktraumen. (German) [On the concept of measure in product spaces]. In Verhandlungen, Internationaler Mathematikerkongress Zürich 1932, volume 2, pages 118–119. Orell Füssli Verlag, Zürich, Switzerland, 1932. LCCN ???? Reprinted in [BMR74].

Ulam:1933:PGI

[Ula33a] Stanisław M. Ulam. Probleme 56: Über gewisse Invarianten der e-Abbildungen. (German) [Problem 56: On certain invariants of etransformations]. Mathematische Annalen, 108(??):311–318, ???? 1933. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Ulam:1933:GZM

[Ula33b] Stanisław M. Ulam. Über gewisse Zerlegungen von Mengen. (German) [On certain decompositions of quantities]. Fundamenta mathematicae, 20(??):221–223, ???? 1933. ISSN 0016-2736 (print), 1730-6329 (electronic).

Ulam:1934:SAM

[Ula34] Stanisław M. Ulam. Über stetige Abbildungen von Mannigfaltigkeiten. (German) [On continuous maps of manifolds]. Annales de la Société Polonaise de Mathématiques, 13(??):141–??, ???? 1934.

Ulam:1938:BTS

[Ula38a] Stanisław M. Ulam. On bounded transformations of space. Preliminary report. Bulletin of the American Mathematical Society, 44(??):195–??, ???? 1938. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1938:DGM

[Ula38b] Stanisław M. Ulam. On the distribution of a general measure in any complete metric separable space. Bulletin of the American Mathematical Society, 44(??):786–??, ???? 1938. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1938:P

[Ula38c] Stanisław M. Ulam. Probleme 74. Fundam. Math., Warszawa, 30 (??):365, 1938.

Ulam:1938:STI

[Ula38d] Stanisław M. Ulam. Set-theoretical invariants of the product operation. Bulletin of the American Mathematical Society, 44(??): 195–??, ???? 1938. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1938:TEF

[Ula38e] Stanisław M. Ulam. Sur les transformations ergodiques. (French) [On ergodic transformations]. Annales de la Société Polonaise de Mathématiques, 17(??):112–??, ???? 1938.

Ulam:1938:TFT

[Ula38f] Stanisław Marcin Ulam. Theory of functions of two real variables. Lectures at Brown University during 1937–38. Brown university, Providence, RI, USA, 1938. 1–38 pp. LCCN QA331 .U6.

Ulam:1939:ITP

[Ula39a] Stanisław M. Ulam. e-isomorphic transformations. Preliminary report. Bulletin of the American Mathematical Society, 45(??):232– ??, ???? 1939. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1939:ATM

[Ula39b] Stanisław M. Ulam. On the abstract theory of measure. Bulletin of the American Mathematical Society, 45(??):83–??, ???? 1939. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1941:MSS

[Ula41a] Stanisław M. Ulam. On measures for subsets of sets of measure zero. Bulletin of the American Mathematical Society, 47(??):702– ??, ???? 1941. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1941:TOP

[Ula41b] Stanisław M. Ulam. Theory of operation of products of sets. I. Preliminary report. Bulletin of the American Mathematical Society, 47(??):702–??, ???? 1941. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1942:GAT

[Ula42] Stanisław M. Ulam. Geometrical approach to the theory of representations of topological groups. preliminary report. Bulletin of the American Mathematical Society, 48(??):44–??, ???? 1942. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1943:WM

[Ula43a] S. M. Ulam. What is measure? American Mathematical Monthly, 50:597–602, 1943. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ulam:1943:EF

[Ula43b] Stanisław M. Ulam. On the equivalence of functions. Bulletin of the American Mathematical Society, 49(??):49–??, ???? 1943. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1943:LCS

[Ula43c] Stanisław M. Ulam. On the length of curves, the surface area and the isoperimetric problem under a general Minkowski metric. Preliminary report. Bulletin of the American Mathematical Society, 49(??):57–??, ???? 1943. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1943:TOP

[Ula43d] Stanisław M. Ulam. Theory of the operation of products of sets. II. Preliminary report. Bulletin of the American Mathematical Society, 49(??):367–368, ???? 1943. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1944:TOP

[Ula44] Stanisław M. Ulam. Theory of the operation of product of sets. III. Preliminary report. Bulletin of the American Mathematical Society, 50(??):60–61, ???? 1944. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Ulam:1946:SB

[Ula46] Stanisław M. Ulam. Stefan Banach, 1892–1945. Bulletin of the American Mathematical Society, 52(7):600–603, ???? 1946. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

URL http://projecteuclid.org/euclid.bams/1183509473.

Ulam:1947:QFP

[Ula47a] Stanisław M. Ulam. On quasi-fixed points for transformations in function spaces. Bulletin of the American Mathematical Society, 53(11):1120, November 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://www.ams.org/

journals/bull/1947-53-11/S0002-9904-1947-08918-7/S0002-
9904-1947-08918-7.pdf.

Ulam:1947:RAM

[Ula47b] Stanisław M. Ulam. Review: Andrzej Mostowski, On Undecidable Propositions in Formalized Systems of Mathematics. Journal of Symbolic Logic, 12(2):60, ???? 1947. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL http://

projecteuclid.org/euclid.jsl/1183395252.

Ulam:1948:SMP

[Ula48] Stanisław M. Ulam. Statistical methods for problems involving equations of the diffusion type, (Monte Carlo). In ????, editor, A.E.C. Information Meeting, Brookhaven National Laboratory, April 26-28, 1948, pages 27–?? ????, ????, 1948. LCCN ???? BNL-17.

Ulam:1949:PPC

[Ula49] Stanisław M. Ulam. Problems in probability and combinatorial analysis. In Proceedings of a conference held on the UCLA campus at the National Bureau of Standards Institute for Numerical Analysis, 29–31 July 1948, page ?? ????, ????, 1949. Cited in [Hur85, page 142] as presented, and using the term “Monte Carlo”, but not published in the proceedings in 1949.

Ulam:1951:MCM

[Ula51] S. Ulam. On the Monte Carlo method. In Anonymous [Ano51], pages 207–212. LCCN QA75 .S9 1949. This publication appears in Science, Computers and People: From the Tree of Mathematics, edited by Mark C. Reynolds and Gian-Carlo Rota. Boston: Birkhaüser, 1986.

Ulam:1952:RPT

[Ula52] S. Ulam. Random processes and transformations. In Proceedings of the International Congress of Mathematicians, Vol. 2, Cambridge, Massachusetts, August 30–September 6, 1950, pages 264– 275. American Mathematical Society, Providence, RI, USA, 1952.

Ulam:1955:HF

[Ula55] Stanisław M. Ulam. Homage to Fermi. Santa Fe New Mexican, ?? (??):??, January 6, 1955.

Ulam:1956:AMC

[Ula56a] S. Ulam. Applications of Monte Carlo methods to tactical games. In Symposium on Monte Carlo methods, University of Florida, 1954, pages 63–?? Wiley, New York, NY, USA, 1956. Abstract only.

Metropolis:1956:PSM

[Ula56b] Stanisław M. Ulam. Applications of Monte Carlo methods to tactical games. In Herbert A. Meyer, editor, Symposium on Monte Carlo methods, University of Florida, March 16–17, 1954, pages 63–?? Wiley, New York, NY, USA, 1956.

Ulam:1956:EBD

[Ula56c] Stanisław M. Ulam. On the ergodic behavior of dynamical systems. Technical Report LA-2055, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1956. Series of lectures on physics of ionized gases.

Ulam:1956:SMP

[Ula56d] Stanisław M. Ulam. Some mathematical problems investigated through computations on electronic machines. American Mathematical Monthly, 63(??):607–??, ???? 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ulam:1957:IMP

[Ula57a] S. M. Ulam. Infinite models in physics. In Applied probability. Proceedings of Symposia in Applied Mathematics, Vol. VII, pages 87–95. McGraw-Hill, New York, NY, USA, 1957. Proceedings of the Seventh Symposium in Applied Mathematics (Brooklyn Polytechnic Institute, April 14–15, 1955).

Ulam:1957:MST

[Ula57b] Stanisław M. Ulam. Marian Smoluchowski and the theory of probabilities in physics. American Journal of Physics, 25(7):475– 481, October 1957. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v25/i7/p475_s1.

Ulam:1957:SNP

[Ula57c] Stanisław M. Ulam. On some new possibilities in the organization and use of computing machines. IBM research report RC-86, IBM, ????, 1957.

Ulam:1957:SB

[Ula57d] Stanisław M. Ulam, editor. The Scottish book. S. Ulam, Los Alamos, NM, USA, 1957. iii + 88 pp. LCCN QA43 .S533 1957; QA43 .U4. An edited translation of a notebook kept at the Scottish Cafe for the Lwow Section of the Société Polonaise de Mathématiques.

Ulam:1958:B

[Ula58a] S. Ulam. Books. Scientific American, 198:127–137, 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ulam:1958:JN

[Ula58b] S. Ulam. John von Neumann, 1903–1957. Bulletin of the American Mathematical Society, 64(3 (part 2)):1–49 + 1, May 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL

http://projecteuclid.org/euclid.bams/1183522369.

Ulam:1958:FUF

[Ula58c] Stanisław M. Ulam. Future uses of future computers. American Chemical Society Abstracts of Papers, 133(??):33–??, ???? 1958.

Ulam:1958:LJN

[Ula58d] Stanisław M. Ulam. The late John von Neumann on computers and the brain. Scientific American, ??(??):127–??, June 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ulam:1958:PEE

[Ula58e] Stanisław M. Ulam. On the possibility of extracting energy from gravitational systems by navigating space vehicles. Technical Report LAMS-2219, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1958.

Ulam:1958:SBU

[Ula58f] Stanisław M. Ulam. Statement before the U.S. House of Representatives. Hearings on astronautics and space exploration. 85th Congress, 2nd session, April 15–May 12, 1958, 1958.

Ulam:1959:RFR

[Ula59a] Stanisław M. Ulam. Review of Funkcje Rzeczywiste [real functions] by Roman Sikorski. Bulletin of the American Mathematical Society, 65(5):305–306, ???? 1959. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://projecteuclid.

org/euclid.bams/1183523312.

Ulam:1959:RRS

[Ula59b] Stanisław M. Ulam. Review of Roman Sikorski, Funkcje Rzeczywiste [Real functions], Volume II. Bulletin of the American Mathematical Society, 67(2):172, ???? 1959. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http://

projecteuclid.org/euclid.bams/1183524071.

Ulam:1960:CMP

[Ula60a] Stanisław M. Ulam. A Collection of Mathematical Problems, volume 8 of Interscience Tracts in Pure and Applied Mathematics. Interscience Publishers, New York, NY, USA, 1960. xiii + 150 pp. LCCN QA43 .U5; QA43 .U36; QA3 .I57. Reprinted in [Ula64c]. Translated into Russian (1964).

Ulam:1960:SBJ

[Ula60b] Stanisław M. Ulam. Statement before the Joint Committee on Atomic Energy. in Frontiers in Atomic Energy Research: Hearings before the Subcommittee on Research and Development of the Joint Committee on Atomic Energy. Eighty-sixth Congress, Second Session, on Frontiers in Atomic Energy Research, March 22–25, 1960, 282–285. Washington, D.C.: U.S. Government Printing Office, 1960, 1960.

Ulam:1961:SSP

[Ula61a] S. M. Ulam. On some statistical properties of dynamical systems. In Lucian M. Le Cam, Jerzy Neyman, and Elizabeth Scott, editors, Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability (University of California, Berkeley, June 20–July 30, 1960), volume 3, pages 315–320. University of California Press, Berkeley, CA, USA, 1961. URL http://projecteuclid. org/euclid.bsmsp/1200512818. Translated into Russian (1963).

Ulam:1961:MCC

[Ula61b] Stanisław M. Ulam. Monte Carlo calculations in problems of mathematical physics. In Edwin F. Beckenbach, editor, Modern Mathematics for the Engineer, 2nd series, chapter 11, pages 261–281. McGraw-Hill, New York, NY, USA, 1961.

Ulam:1961:SMS

[Ula61c] Stanisław M. Ulam. On a statistical method of solving multiplication and diffusion problems. In Monsanto Chemical Company Meeting, Clinton National Laboratory, Oct. 13–15, 1947, volume Mon-411(14) of Monsanto Chemical Company Abstracts of Papers, page ?? ????, ????, 1961. LCCN ????

Ulam:1962:SMB

[Ula62a] S. Ulam. Stability of many-body computations. In Garrett Birkhoff, Richard Bellman, and C. C. Lin, editors, Hydrodynamic Instability, volume 13 of Symposia in Applied Mathematics, pages 247–258. American Mathematical Society, Providence, RI, USA, 1962.

Ulam:1962:SMP

[Ula62b] S. M. Ulam. On some mathematical problems connected with patterns of growth of figures. Proceedings of symposia in applied mathematics, 14:215–224, 1962. CODEN PSYMA7. ISSN 0160-7634. An early study on pattern development in cellular automata by the man who suggested CA’s to von Neumann. Reprinted in [Bur70].

Ulam:1962:ECS

[Ula62c] Stanisław M. Ulam. Electronic computers and scientific research. In Carl F. J. Overhage, editor, The Age of Electronics, pages 95– 108. McGraw-Hill, New York, NY, USA, 1962. LCCN ???? Also in Computers and Automation, August 1963 and September 1963.

Ulam:1962:MCC

[Ula62d] Stanislaw M. Ulam. Monte Carlo calculations in problems of mathematical physics. In Beckenbach [Bec62], page ?? LCCN QA401 .B4. With an introduction by Magnus R. Hestenes.

Ulam:1962:OP

[Ula62e] Stanisław M. Ulam. An open problem. In Anonymous [Ano62], page 223. LCCN QA269.P75.

Ulam:1963:NLT

[Ula63a] Stanisław M. Ulam. Communication to the U.S. Senate Committee on Foreign Relations. in Nuclear Test Ban Treaty: Hearings before the Committee on Foreign Relations, United States Senate. Eightyeighth Congress, First Session, on Executive M, August 12–15, 19– 23, 26–27, 1963. Washington, DC: US Goverment Printing Office, 1963.

Ulam:1963:SPD

[Ula63b] Stanisław M. Ulam. Some problems of a dilettante. problems 110, 111, and 112. In S. Chowla and B. Jones, editors, Proceedings of the 1963 Number Theory Conference (University of Colorado, Boulder, Colorado, August, 1963), pages 114–115. ????, ????, 1963. LCCN ????

Ulam:1963:SPC

[Ula63c] Stanisław M. Ulam. Some properties of certain non-linear transformations. In Stefan Drobot, editor, Mathematical Models in Physical Sciences: Proceedings of the Conference at the University of Notre Dame, 1962, pages 85–95. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1963. LCCN ????

Ulam:1964:CAI

[Ula64a] S. Ulam. Combinatorial analysis in infinite sets and some physical theories. SIAM Review, 6(4):343–355, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ulam:1964:C

[Ula64b] Stanisław M. Ulam. Computers. Scientific American, 211:202– 216, 1964. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ulam:1964:PMM

[Ula64c] Stanisław M. Ulam. Problems in modern mathematics. Science editions. Wiley, New York, NY, USA, 1964. xvii + 150 pp. LCCN QA43 .U5 1964.

Ulam:1965:OP

[Ula65a] Stanisław M. Ulam. The Orion project. Nuclear News, ??(??): 25–27, January 1965. CODEN NUNWA8. ISSN 0029-5574.

Ulam:1965:RDS

[Ula65b] Stanisław M. Ulam. Recursive definitions of static changing patterns. In ????, editor, Biomathematics and Computer Science in the Life Sciences, volume 9 of Monograph of Proceedings of Third Annual Symposium, Houston, Texas, (1965), page ?? ????, ????, 1965. LCCN ????

Ulam:1966:CSS

[Ula66a] Stanisław M. Ulam. Collapse of stellar systems. In ????, editor, Proceedings of the 25th International Astronomical Union Symposium (Thessaloniki, Greece, August 16–22, 1964), pages 76–77. International Astronomical Union, ????, 1966. LCCN ????

Ulam:1966:HFM

[Ula66b] Stanisław M. Ulam. How to formulate mathematically problems of the rate of evolution? In Paul S. Moorhead and Martin M. Kaplan, editors, Proceedings of the Symposium on Mathematical Challenges to the Neo-Darwinian Interpretation of Evolution (New York, April 5–8, 1961), volume 5 of Wistar Institute Monograph, pages 21– 33. American Mathematical Society, Providence, RI, USA, 1966. LCCN ???? A. Liss, 1985.

Ulam:1966:MC

[Ula66c] Stanisław M. Ulam. La machine creatrice. In Caillois et al. [CU^{+}66], pages 31–42. LCCN ????

Ulam:1966:TD

[Ula66d] Stanisław M. Ulam. Thermonuclear devices. In R. E. Marshak and J. Warren Blaker, editors, Perspectives in Modern Physics: Essays in Honor of Hans A. Bethe, pages 593–601. Interscience Publishers, New York, NY, USA, 1966. LCCN ????

Ulam:1967:HFM

[Ula67a] S. M. Ulam. How to formulate mathematically problems of rate of evolution? The Wistar Institute symposium monograph, 5(??): 21–33, ???? 1967. ISSN 0084-1013.

Ulam:1967:EAM

[Ula67b] Stanisław M. Ulam. An education in applied math. SIAM Review, 9(??):343–344, 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Proceedings of May 24–27, 1966 SIAM Conference (Aspen, Colorado).

Ulam:1967:GFS

[Ula67c] Stanisław M. Ulam. On general formulations of simulation and model construction. In George Shapiro and Milton Rogers, editors, Prospects for Simulation and Simulators of Dynamic Systems, pages 3–8. Spartan Books, New York, NY, USA, 1967. LCCN ????

Ulam:1968:CCB

[Ula68a] S. M. Ulam. Computations on certain binary branching processes. In R. F. Churchhouse and J.-C. Herz, editors, Computers in Mathematical Research, pages 168–171. North-Holland Publishing Co., Amsterdam, The Netherlands, 1968.

Ulam:1968:NSS

[Ula68b] Stanisław M. Ulam. Numerical studies of star systems. In Colloque sur le Problème des N Corps, pages 265–267. Editions du Centre National de la Recherche, ????, 1968. LCCN ????

Ulam:1968:PIS

[Ula68c] Stanisław M. Ulam. Philosophical implications of some recent scientific discoveries. In Science, Philosophy and Religion. Proceedings of the Symposium held at Kirtland Air Force Laboratory, Albuquerque, New Mexico, pages 44–48. ????, ????, 1968. LCCN ????

Ulam:1969:RSC

[Ula69a] S. Ulam. Reminiscences from the Scottish Café. Wiadom. Mat. (2), 12:49–58 + 2, 1969. ISSN 0373-8302.

Ulam:1969:AM

[Ula69b] Stanisław M. Ulam. The applicability of mathematics. In The Mathematical Sciences: A Collection of Essays, pages 1–6. MIT Press, Cambridge, MA, USA, 1969. LCCN ???? Edited by the National Research Council’s Committee on Support of Research in the Mathematical Sciences.

Ulam:1969:WKS

[Ula69c] Stanisław M. Ulam. Wspomnienia kawiarni szkockiej (reminiscences of the Scottish Café). Wiadomosci Matematyczne, 12(??): 49–58, ???? 1969.

Ulam:1970:F

[Ula70] Stanisław M. Ulam. Foreword. In My World Line: An Informal Autobiography [Gam70b], page ?? LCCN ???? URL http://adsabs.

harvard.edu/abs/1970mwla.book.....G. Foreword by Stanisław

M. Ulam.

Ulam:1971:GPI

[Ula71a] Stanisław M. Ulam. Generalizations of product isomorphisms. In Capobianco et al. [CFK71], pages 215–?? ISBN 3-540-05386-7 (Berlin), 0-387-05386-7 (New York). LCCN ????

Ulam:1971:TBU

[Ula71b] Stanisław M. Ulam. Testimony before the United States District Court, District of Minnesota, Minneapolis, Minnesota, September 17, 1971, in the case of Honeywell Incorporated versus Sperry Rand Corporation, 7342-438, 1971.

Ulam:1972:SCP

[Ula72a] S. M. Ulam. Some combinatorial problems studied experimentally on computing machines. In S. K. Zaremba, editor, Applications of Number Theory to Numerical Analysis (Proceedings of a Symposium held at Université de Montréal, Montréal, Québec, 1971), pages 1–10. Academic Press, New York, NY, USA, 1972.

Ulam:1972:SIP

[Ula72b] S. M. Ulam. Some ideas and prospects in biomathematics. In M. F. Morales, editor, Annual Review of Biophysics and Bioengineering, volume 1, pages 277–291. Annual reviews Inc., Palo Alto, CA, 1972.

Ulam:1972:GM

[Ula72c] Stanisław M. Ulam. Gamow—and mathematics. In Reines [Rei72], pages 27–29. ISBN 0-87081-025-1. LCCN QC780 .C65.

Ulam:1972:ISS

[Ula73] Stanisław M. Ulam. Ideas of space and space-time. Rehovot Magazine, ??(??):29–33, Winter 1972/1973.

Ulam:1974:I

[Ula74a] Stanisław M. Ulam. Infinities. In The heritage of Copernicus: theories “pleasing to the mind” [Ney74], pages 378–392. ISBN 0-262-14021-7. LCCN Q126.8 .N48.

Ulam:1974:NRO

[Ula74b] Stanisław M. Ulam. New rules and old games. Outlook, ??(??): 32–33, Spring 1974.

Ulam:1975:AKD

[Ula75] Stanisław M. Ulam. Arthur Koestler et le défi du hazard: Entretien avec Stan Ulam de Pierre Debray-Ritzen. (French) [Arthur Koestler and the challenge of chance: Interview with Stan Ulam by Pierre Debray-Ritzen]. In Arthur Koestler, Cahiers de l’Herne, pages 428– 432. Edition de L’Herne, Paris, France, 1975. ISBN ???? LCCN ????

Ulam:1976:AM

[Ula76a] Stanisław M. Ulam. Adventures of a mathematician. Charles Scribner’s Sons, New York, NY, USA, 1976. ISBN 0-684-14391-7. xi + 317 + 25 pp. LCCN QA29.U4 A33; QA29.U4 A33 1976; QA29.U4 A331. Paperback editions published in 1977 and 1983. Translated into Japanese (1979).

Ulam:1976:PM

[Ula76b] Stanisław M. Ulam. Physics for mathematicians. In Huang [Hua76], pages 113–1212. ISBN 0-88318-127-4. LCCN QC28 .P48.

Ulam:1977:SB

[Ula77] Stanisław M. Ulam. The Scottish Book. LASL Monograph LA-6832, Los Alamos Scientific Laboratory, Los Alamos, NM, USA,

June 1977. vi + 56 pp. URL https://www.osti.gov/servlets/

purl/7218289. Reprint of [Mau81].

Ulam:1978:BI

[Ula78a] Stanisław M. Ulam. Banach i inni. Kultura, 10(??):??, Sierpnia 6, 1978. Translated into Polish by Jerzy Jaruzelski.

Ulam:1978:NKS

[Ula78b] Stanisław M. Ulam. Narodziny “ksiegi szkockiej”. Kultura, 10(??): ??, Sierpnia 13, 1978. Translated into Polish by Jerzy Jaruzelski.

Ulam:1978:PM

[Ula78c] Stanisław M. Ulam. Przygody matematyka. Kultura, 9:??, Lipca 30, 1978. Translated into Polish by Jerzy Jaruzelski.

Ulam:1978:RAM

[Ula78d] Stanisław M. Ulam. The role of abstract mathematical ideas in possible conceptual advances in natural sciences, more specifically biology. In Dominique Akl, Moshe Flato, and Daniel Sternheimer, editors, Proceedings of International Colloquium on the Role of Mathematical Physics in the Development of Science (College de France, Paris, June 13–15, 1977), pages 12–25. UNESCO, ????, 1978. ISBN ???? LCCN ????

Ulam:1980:OPP

[Ula80a] S. Ulam. On the operations of pair production, transmutations, and generalized random walk. Advances in Applied Mathematics, 1(1):7–21, 1980. ISSN 0196-8858 (print), 1090-2074 (electronic).

Ulam:1980:F

[Ula80b] Stanisław M. Ulam. Foreword. In A half century of Polish mathematics: remembrances and reflections [Kur80], page ?? ISBN 83-01-00819-9. LCCN QA27.P67 K7813. Preface by Stanisław M. Ulam.

Ulam:1980:MPL

[Ula80c] Stanisław M. Ulam. A mathematical physicist looks at computing. Rehovot Magazine, 9(1):47–50, 1980.

Ulam:1980:NIM

[Ula80d] Stanisław Marcin Ulam. Von Neumann: The interaction of mathematics and computing. In Metropolis et al. [MHR80], pages 93–99. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Ulam:1981:FAM

[Ula81a] S. M. Ulam. Future applications of mathematics in the natural sciences. In Tarwater et al. [TWHM81], pages 101–114.

Ulam:1981:KK

[Ula81b] Stanisław M. Ulam. Kazimierz Kuratowski, 1896–1980. Polish Review, 26(??):62–66, 1981.

Ulam:1981:NAC

[Ula81c] Stanisław M. Ulam. On the notion of analogy and complexity in some constructive mathematical schemata. Technical Report LA-9065, Los Alamos National Laboratory, Los Alamos, NM, USA, 1981. Also in [Rot86b].

Ulam:1982:BJN

[Ula82a] S. Ulam. Biographies: John von Neumann, 1903–1957. Annals of the History of Computing, 4(2):157–181, April/June 1982. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1982/pdf/a2157.pdf;   http://www.computer.org/
annals/an1982/a2157abs.htm.

Ulam:1982:JN

[Ula82b] S. Ulam. John von Neumann, 1903–1957. Annals of the History of Computing, 4(2):157–181, 1982. CODEN AHCOE5. ISSN 0164-1239. Reprint.

Ulam:1982:AHS

[Ula82c] Stanisław M. Ulam. An anecdotal history of the Scottish Book. In Mauldin [Mau81], page ?? ISBN 3-7643-3045-7. LCCN QA43 .S39. First edition 1957 and 1977 by Stanisław Ulam and distributed by the Los Alamos Laboratory. Contains mathematical problems, collected in a note-book by Polish mathematicians in the Scottish coffee house in Lwów from 1935 to 1941. Includes selected papers from the Scottish Book Conference, held at North Texas State University in May 1979.

Ulam:1982:I

[Ula82d] Stanisław M. Ulam. Introduction. In Rassias and Rassias [RR82], page ?? ISBN 0-444-86161-0. LCCN QB461 .S374 1982.

Ulam:1982:SAM

[Ula82e] Stanisław M. Ulam. Speculations about the mechanism of recognition and discrimination. Technical Report LAUR 82-62, Los Alamos National Laboratory, Los Alamos, NM, USA, 1982.

Ulam:1982:TIM

[Ula82f] Stanisław M. Ulam. Transformations, iterations and mixing flows. In Bednarek and Cesari [BC82], pages 419–426. ISBN 0-12-084720-5, 0-12-083750-1. LCCN QA614.8 .D93 1982.

Ulam:1983:KKW

[Ula83] Stanisław M. Ulam. Kazimierz Kuratowski, wspomnienia. (Polish) [Kazimierz Kuratowski: A reminiscence]. Wiadomosci Matematyczne, ??(??):??, ???? 1983. Translated into Polish by R. Engelking. Also in Kazimierz Kuratowski, Selected Papers, Polish Academy of Sciences, K. Borsuk, editor, PWN, Warsaw, 1988.

Ulam:1985:SSP

[Ula85] S. Ulam. Speculations on some possible mathematical frameworks for the foundations of certain physical theories. Letters in Mathematical Physics, 10(2–3):101–106, 1985. CODEN LMPDHQ. ISSN 0377-9017 (print), 1573-0530 (electronic).

Ulam:1986:NAC

[Ula86] S. M. Ulam. On the notion of analogy and complexity in some constructive mathematical schemata. In Rota [Rot86b], pages 35– 45. ISBN 0-12-598543-6. LCCN QA273 .P7954 1986.

Ulam:1987:CR

[Ula87a] Françoise Ulam. Conversations with Rota. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):300–312, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-25.pdf.

Ulam:1987:SUE

[Ula87b] Françoise Ulam. Stan Ulam: Esquisse. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):4–7, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-
bin/getfile?15-01.pdf.

Ulam:1987:RBA

[Ula87c] S. M. Ulam. Reflections on the brain’s attempts to understand itself. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):283–287, 1987. CODEN LASCDI. ISSN 0273-

7116. URL http://library.lanl.gov/cgi-bin/getfile?15-20.

pdf. A Gamow Memorial Lecture delivered at the University of Colorado. Boulder, on October 5, 1982. With a contribution by William A. Beyer: “An Ulam Disance”.

Ulam:1987:SUV

[Ula87d] S. M. Ulam. Stan Ulam: Vita — excerpts from adventures of a mathematician. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):8–22, 1987. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?

15-02.pdf. Stanisław Ulam 1909–1984.

Ulam:1987:SRT

[Ula87e] S. M. Ulam. Sub rosa — A trialogue. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):295–299, 1987. CODEN

LASCDI. ISSN 0273-7116. URL http://library.lanl.gov/cgi-

bin/getfile?15-24.pdf. Stanisław Ulam 1909–1984.

Ulam:1991:AM

[Ula91a] Stanisław M. Ulam. Adventures of a mathematician. University of California Press, Berkeley, CA, USA, 1991. ISBN 0-520-07154-9. xxxiv + 329 + 22 pp. LCCN QA29.U4 A33 1991eb; QA29.U4; Internet.

Ulam:1991:A

[Ula91b] Stanisław M. Ulam. Los Alamos. In Ferris and Fadiman [FF91], pages 706–729. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Ulam:1994:RAM

[Ula94] Françoise Ulam. Reminiscences about S. M. Ulam. In Themistocles M. Rassias and Józef Tabor, editors, Stability of mappings of Hyers–Ulam type, Hadronic Press Collect. Orig. Artic., pages 3–5. Hadronic Press, Palm Harbor, FL, USA, 1994. ISBN 0-911767-82-7, 0-911767-64-9 (paperback). LCCN QH431 S73x, 1994.

Ulam:1997:MTG

[Ula97] Stanisław Ulam. On measure theory in general set theory (doctoral dissertation). Wiadom. Mat., 33:155–168, 1997. ISSN 0373-8302. Reprint of the 1933 original.

Ulam:2013:MCC

[Ula13] Stanislaw M. Ulam. Monte Carlo calculations in problems of mathematical physics. In Beckenbach [Bec13], page ?? ISBN 0-486-31612-2 (e-book), 0-486-49747-X (paperback), 1-306-32668-0 (ebook). LCCN QA401 .B39 2013. With an introduction by Magnus R. Hestenes. Reprint of [Bec62].

Ulam:1953:HSP

[UP53] Stanisław M. Ulam and J. Pasta. Heuristic studies in problems of mathematical physics on high speed computing machines. Technical Report LA-1557, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1953.

Ulam:1986:GMP

[URR86a] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Gamow and mathematics: Personal reminiscences. In Science, computers, and people: from the tree of mathematics [URR86d], pages 231–240. ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:JN

[URR86b] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. John von Neumann, 1903–1957. In Science, computers, and people: from the tree of mathematics [URR86d], chapter 16, pages 169–214.

ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:JNC

[URR86c] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. John von Neumann on computers and the brain. In Science, computers, and people: from the tree of mathematics [URR86d], chapter 18, pages 223–230. ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:SCP

[URR86d] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Science, computers, and people: from the tree of mathematics. Birkhäuser Boston Inc., Cambridge, MA, USA, 1986. ISBN 0-8176-3276-X. xxii + 264 pp. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1986:NIM

[URR86e] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Von Neumann: The interaction of mathematics and computing. In Science, computers, and people: from the tree of mathematics [URR86d], chapter 17, pages 215–222. ISBN 0-8176-3276-X. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ulam:1931:PMM

[US31] Stanisław M. Ulam and J. Schreier. Sur une propriété de la mesure de M. Lebesgue. (French) [On a property of the measure of M. Lebesgue]. Comptes Rendus de l’Academie des Sciences, 192(??): 539–542, ???? 1931.

Ulam:1933:GPS

[US33] Stanisław M. Ulam and J. Schreier. Sur le groupe des permutations de la suite des nombres naturels. (French) [On the permutation group of sequences of natural numbers]. Comptes rendus de l’Académie des sciences, Paris, 197(??):737–738, ?? 1933.

Ulam:1971:SEA

[US71] S. Ulam and R. Schrandt. Some elementary attempts at numerical modeling of problems concerning rates of evolutionary processes. Technical Report LAMS-4573, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1971. See [US86].

Ulam:1986:SEA

[US86] S. Ulam and R. Schrandt. Some elementary attempts at numerical modeling of problems concerning rates of evolutionary processes. Physica D, Nonlinear phenomena, 22(1–3):4–12, 1986. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic). Evolution, games and learning (Los Alamos, NM, 1985).

Ulam:1947:CSD

[UvN47] S. M. Ulam and John von Neumann. On combinations of stochastic and deterministic processes. Bulletin of the American Mathematical Society, 53(11):1120, November 1947. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

URL http://www.ams.org/journals/bull/1947-53-11/S0002-
9904-1947-08918-7/S0002-9904-1947-08918-7.pdf. The ab-

stract notes “… starting with almost every x_{1} (in the sense of Lebesgue measure) and iterating the function f(x) = 4 · (1 − x) one obtains a sequence of numbers on (0, 1) with a computable algebraic distribution. By playing suitable games with numbers ‘drawn’ in this fashion, one can obtain various other distributions, either given explicitly or satisfying given differential or integral equations.”.

Ulam:1964:PAP

[UW64] Stanisław M. Ulam and W. E. Walden. Possibility of an accelerated process of collapse of stars in a very dense centre of a cluster or a galaxy. Nature, 201(4925):1202, March 21, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/2011202a0.

vonNeumann:1966:TSR

[vNB66] John von Neumann and Arthur W. (Arthur Walter) Burks, editors. Theory of Self-Reproducing Automata. University of Illinois Press, Urbana, IL, USA, 1966. xix + 388 pp. LCCN QA267 .V55. Von Neumann’s work on self-reproducing automata, completed and edited after his death by Arthur Burks. Also includes transcripts of von Neumann’s 1949 University of Illinois lectures on the “Theory and Organization of Complicated Automata”. See [Bur70].

vonNeumann:1945:RET

[vNU45] John von Neumann and S. M. Ulam. Random ergodic theorems. Bulletin of the American Mathematical Society, 51(??):660–??, ???? 1945. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

vonNeumann:1947:GHS

[vNU47] John von Neumann and S. Ulam. On the group of homeomorphisms of the surface of the sphere. Bulletin of the American Mathematical Society, 53(??):506–??, ???? 1947. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Waterman:2006:SUC

[Wat06] M. S. Waterman. Stan Ulam and computational biology. In A. Apostolico, C. Guerra, S. Istrail, P. Pevzner, and M. Waterman, editors, Research in Computational Molecular Biology, Proceedings, volume 3909, page 159. Springer-Verlag Inc., New York, NY, USA, 2006. ISBN 3-540-33295-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Weissert:1997:GSD

[Wei97] Thomas P. Weissert. The Genesis of Simulation in Dynamics: Pursuing the Fermi–Pasta–Ulam Problem. Springer-Verlag Inc., New York, NY, USA, 1997. ISBN 0-387-98236-1, 0-387-98237-X. xiii + 176 pp. LCCN QC133 .W45 1997.

Weisstein:2006:US

[Wei06] Eric W. Weisstein. Ulam sequence, 2006. URL http://mathworld.

wolfram.com/UlamSequence.html. From MathWorld — a Wol-

fram Web resource.

Wellerstein:2012:ITU

[Wel12] Alex Wellerstein. The infamous Teller–Ulam report (1951). Web

site., February 22, 2012. URL http://blog.nuclearsecrecy.
com/2012/02/22/weekly-document-15-the-infamous-teller-
ulam-report-1951/.

Wellerstein:2017:SSH

[WG17] Alex Wellerstein and Edward Geist. The secret of the Soviet hydrogen bomb. Physics Today, 70(4):40–47, April 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wilson:1976:BMB

[Wil76] Jane Wilson. Books: Math beneath the ivory tower: Adventures of a Mathematician, by S. M. Ulam. Bulletin of the Atomic Scientists, 32(9):45–46, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1987:MWT

[Yor87] Herbert F. (Herbert Frank) York. Making weapons, talking peace: a physicist’s odyssey from Hiroshima to Geneva. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-04338-0. xiv + 359 + 8 pp. LCCN JX1974.7 .Y575 1987. US$22.95.

Zielinski:1977:LPP

[Zie77] Ed Zielinski. Letter: Polish and proud. Bulletin of the Atomic Scientists, 33(5):7, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zorpette:2024:QRG

[Zor24] Glenn Zorpette. 5 questions: Richard Garwin: The designer of the first hydrogen bomb on creating “the sausage”. IEEE Spectrum, 61(9):22–23, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).